### v1

# On Exact Space-Depth Trade-Offs in Multi-Controlled Toffoli Decomposition

Abstract
In this paper, we consider the optimized implementation of Multi Controlled Toffoli (MCT) using the Clifford $+$ T gate sets. While there are several recent works in this direction, here we explicitly quantify the trade-off (with concrete formulae) between the Toffoli depth (this means the depth using the classical 2-controlled Toffoli) of the $n$-controlled Toffoli (hereform we will tell $n$-MCT) and the number of clean ancilla qubits. Additionally, we achieve a reduced Toffoli depth (and consequently, T-depth), which is an extension of the technique introduced by Khattar et al. (2024). In terms of a negative result, we first show that using such conditionally clean ancilla techniques, Toffoli depth can never achieve exactly $\lceil\log_{2}n\rceil$, though it remains of the same order. This highlights the limitation of the techniques exploiting conditionally clean ancilla [Nie et al., 2024, Khattar et al., 2024]. Then we prove that, in a more general setup, the T-Depth in the Clifford + T decomposition, via Toffoli gates, is lower bounded by $\lceil\log_{2}n\rceil$, and this bound is achieved following the complete binary tree structure. Since the ($2$-controlled) Toffoli gate can further be decomposed using Clifford $+$ T, various methodologies are explored too in this regard for trade-off related implications.
I Introduction
Quantum gates are the fundamental building blocks of quantum circuits. Unlike classical gates, quantum gates are inherently reversible and are mathematically represented by unitary matrices. The doubly-controlled X-gate, commonly known as the Toffoli gate, is among the most significant quantum gates, with critical applications in arithmetic operations, reversible computing, and oracle constructions. However, the Toffoli gate’s high resource demands, particularly in terms of T-Count, T-Depth, and ancilla qubits, can greatly influence the efficiency of fault-tolerant quantum circuits. Consequently, optimizing its implementation is crucial for reducing computational overhead, minimizing error rates, and enhancing scalability, thereby making it indispensable for practical large-scale quantum computations.
An $n$-controlled Toffoli is an $(n+1)$-qubit gate, where $n$ many qubits control the outcome of a single target qubit. Like ($2$-controlled) Toffoli, Multi Controlled Toffoli (MCT) also has a huge applications in quantum arithmetic, error correction, and implementation of quantum algorithms. For example, in Grover’s Search, implementing the $n$-bit AND function requires an $n$-controlled Toffoli ($n$-MCT) gate. As MCT cannot be implemented in its original form (so far), it needs to be decomposed into Clifford + Toffoli, and eventually to the Clifford + T gate set. Throughout the paper, By Toffoli, we refer to $n$-controlled Toffoli with $n=2$.
It is well known that decomposing/ designing larger circuits with smaller components requires additional qubits, known as ancilla qubits, to reduce depth. The important task for optimizing the number of gates, the depth and the number of additional qubits has received serious attention for a long time and one may refer to the seminal work of Moore and Nilsson moore01 two decades back. A more recent and comprehensive discussion is available in jiang20 and we like to quote from that work “Can we characterize the relationship between the number of ancilla and the possible optimal depth?” The paper jiang20 is concerned with CNOT, whereas in this paper we concentrate on Toffoli. However, the fundamental motivation of quantum circuit synthesis revolves around this very question.
Over the past decades, numerous efforts have been made to reduce the resource requirements for the optimized implementation of multi-controlled Toffoli (MCT) gates miller11 ; saeedi13 ; maslov16 ; baker19 ; paler22 ; bala22 ; claudon24 . More recent works nie24 ; khattar24 have introduced the conditionally clean ancilla technique, which significantly reduces both the Toffoli depth and ancilla count in the MCT decomposition.
Before proceeding further let us explain the concept of conditionally clean ancilla that will be repeatedly referred in this work. Generally, ancilla qubits are of two types, namely the clean ancilla, which is initialized to |$\ket{0}$ at the beginning of the circuit, and the dirty ancilla, which is initialized to some unknown quantum state at the beginning. The standard practice is to re-initialize the ancilla qubit to its initial state at the end of the computation. Conditionally clean ancilla nie24 ; khattar24 are the set of working (control) qubits, derived form the existing ones, that are treated as the clean ancilla based on certain assumptions, and re-initialized at the end of the computation. Although, earlier works nie24 ; khattar24 did consider the dirty ancilla, in this paper, we are focusing on the conditionally clean ancilla derived from the clean ones only. Very recently, the conditionally clean ancilla technique has gained significant attention, including its applications in quantum adders remaud25 .
The resource optimization metrics under consideration are the Toffoli count , Toffoli depth and Ancilla count. The Toffoli count and Toffoli depth are further refined with the T-Count and T-Depth as one may refer to Table 1.
In this paper, we explore how the Toffoli depth (and consequently, T-Depth) can be reduced by increasing the number of clean ancilla qubits, utilizing the concept of conditionally clean ancilla. We also establish the limitation of this method in terms of the lower bound of the Toffoli depth. Additionally, we show that, in a more general setting, the Toffoli depth (which can be further reduced to T-Depth) of an $n$-MCT decomposition cannot be reduced beyond $\lceil\log_{2}n\rceil$. The section-wise contributions of this paper are outlined as follows.
I.1 Organization and Contributions
In Section II, we proceed with the preliminaries. Section III is a warm-up section where we first present the existing results on the Clifford $+$ T decomposition of ($2$-controlled) Toffoli and summarize the results in terms of T-Count, T-Depth, and ancilla requirements in Table 1. Additionally, we also explain the recent developments of MCT decompositions describing the existing best results in terms of Toffoli count, Toffoli depth, and ancilla, which can be subsequently decomposed into T-Count and T-Depth, as summarized in Table 2. Given that there are several developments in very recent times, this section provide a holistic view of the existing results. Based on this, we present our contributions.
In section IV, we explore the trade-off between the (clean) ancilla and the Toffoli depth using the existing techniques related to conditionally clean ones. In Section IV.1, we take a different look in viewing the MCT circuit decomposition using conditionally clean ancilla technique of khattar24 , and enumerate their exact Toffoli count. Following the trade-off, we show the reduction in Toffoli depth, and therefore in T-Depth (Construction 1, Example 1) compared to the recent work of Khattar and Gidney khattar24 , by introducing additional clean ancilla into the circuit, while keeping the Toffoli count constant. These results are shown in Section IV.2. Then, in Section IV.3, we identify the limitation of this technique in Theorem 3, showing that this direction cannot reduce the Toffoli depth to $\lceil\log_{2}n\rceil$, though it is of order $\mathcal{O}(\log_{2}n)$.
In Section V, we prove within a general framework that the exact Toffoli depth in the ($2$-controlled) Toffoli decomposition of an $n$-MCT cannot be less than $\lceil\log_{2}n\rceil$, regardless of the number of ancilla qubits used. In fact, this lower bound can be achieved for T-Depth as well, by the construction of jaques19 . That is, using the technique of jaques19 , we can obtain an $n$-MCT by further decomposing into the Clifford + T gate set with an exact T-Depth of $\lceil\log_{2}n\rceil$ too, using $22n-$ 2 ancilla, and a T-Count of $4(n-1)$. Moreover, using the logical-AND circuit by Gidney gidney18 , the ancilla count can be reduced to $n-2$ keeping the T-Count constant. However, the exact T-Depth in case of gidney18 becomes one more, i.e., $\lceil\log_{2}n\rceil+1$. To highlight our contribution, we are looking at the exact counts and depth instead of their complexity order.
Section VI concludes the paper with a brief summary of our work and outlines the open problems in the related domain.
II Preliminaries
In this section, we briefly proceed with the preliminaries. A qubit is the fundamental unit of quantum information represented as |$\ket{\psi}=\alpha\ket{0}+\beta\ket{1}$, where $\alpha,\beta\in\mathbb{C}$ and $|\alpha|^{2}+|\beta|^{2}=1$. The basis states can also be written as column matrices, as follows.
|$\ket{0}=\begin{pmatrix}1\\ 0\end{pmatrix},\ket{1}=\begin{pmatrix}0\\ 1\end{pmatrix},\ket{\psi}=\alpha\ket{0}+\beta\ket{1}=\begin{pmatrix}\alpha\\ \beta\end{pmatrix}.$
Upon measurement, the qubit collapses to one of the basis states, |$\ket{0}$ with probability $|\alpha|^{2}$ or |$\ket{1}$ with probability $|\beta|^{2}$. Similarly, an $n$-qubit state is described by $2^{n}$ parameters as |$\ket{\psi_{n}}=\sum_{{x}\in\{0,1\}^{n}}\alpha_{{x}}\ket{{x}}$ with the normalization condition $\sum_{x\in\{0,1\}^{n}}|\alpha_{{x}}|^{2}=1$.
The Quantum gates, unlike the classical ones, are inherently reversible and represented by unitary matrices $U$, with inverses denoted by $U^{\dagger}$.
$U^{\dagger}\left(U\ket{\psi}\right)=U\left(U^{\dagger}\ket{\psi}\right)=\ket{\psi}.$
In classical computing, (2-input 1-output) NAND and NOR gates are considered universal as they can be used to construct any classical logic circuit. In contrast, quantum computing involves infinitely many quantum gates, including both single-qubit and multi-qubit ones, making it challenging to define a universal description. However, there exist certain gate sets that can approximate any unitary transformation on a quantum computer to an arbitrary degree of accuracy, known as the universal gate sets. The most common examples of quantum universal gate sets include the Clifford + T gate set, rotation gates combined with the CNOT gate set, etc.
The Clifford group is generated by three gates: Hadamard (H), phase (S), and CNOT. This set is minimal, as removing any one gate would result in the inability to implement some Clifford operations. Since all the Pauli matrices can be derived from the phase and Hadamard gates (Equation 2), each Pauli gate is also an element of the Clifford group.
[…]
However, the Clifford gates alone do not constitute a universal set of quantum gates because certain gates. For example, the T $=\sqrt{\text{S}}$ gate, cannot be arbitrarily approximated using only Clifford operations. Therefore, the Clifford group, when augmented with the T gate, forms a universal quantum gate set for quantum computation.
The Toffoli gate is a three-qubit quantum gate, where the first two serve as control qubits, and the third one is the target. The gate flips the target qubit if and only if both control qubits are in the |$\ket{1}$ state, described as
Toffoli :|$\text{Toffoli }:\ket{x,y,z}\rightarrow\ket{x,y,z\oplus xy}.$
[…]
Figure 1: Equivalence between multi-controlled Pauli gates.
In quantum computing, the Multi-Controlled Toffoli (MCT) gates play a crucial role in the design of complex quantum algorithms, including error correction codes and arithmetic operations. Despite its utility, the implementation of MCT gates exploiting the Clifford + T gate set, requires careful decomposition to optimize resource usage, such as minimizing the T-Count, T-Depth, and the number of ancillary qubits. These optimizations are essential for practical quantum computation, where resource efficiency is a critical consideration. In this regard, we present a comprehensive overview of the existing benchmarks for Clifford + T decompositions of single and multi-controlled Toffoli gates in the following section.
Let us now describe the idea of conditionally clean ancilla from khattar24 , as described in Figure 2. Given a clean ancilla, a Toffoli gate is implemented targeting the clean ancilla. If the clean ancilla is reversed, it means both the control qubits were |$\ket{1}$; thus, applying X gates on the control qubits will change them to |$\ket{0}$ and consequently can be used as conditionally clean ancilla in the next rounds. Similarly, in the following round, additional Toffoli gates are implemented targeting these conditionally clean ancilla and applying the X gate on the control qubits will again make them conditionally clean for the next round, and the process continues until we exhaust all the control qubits. Once the information from all the control qubits is accumulated in a few qubits, another (smaller) MCT gate is implemented with these qubits along with the ancilla as the control qubits and the original target as its target, so that if the ancilla was not modified in the first step, the target will also not be modified.
Figure 2: Understanding the conditionally clean ancilla technique from khattar24 .
Let us now analyze Figure 2 by distributing the possible scenarios in two mutually exclusive and exhaustive cases:
Case I: $C_{0}=C_{1}=C_{2}=C_{3}=C_{4}=\ket{1}$. Then, after Step I, ancilla = |$\ket{1}$. Applying X-gate on $C_{0},C_{1}$ will make, $C_{0}=C_{1}=\ket{0}$, thus can be used as ancilla for Step II. Since, $C_{2}=C_{3}=\ket{1}$, the Toffoli gate will make, $C_{0}=\ket{1}$. Similarly, $C_{0}=\ket{1}$, and $C_{4}=\ket{1}$, imply that $C_{1}=\ket{1}$. Finally, $C_{1}=\ket{1}$, and ancilla =|$=\ket{1}$, will reverse the target qubit.
Case II: At least one of $C_{i}=\ket{0}$, for some $i\in[0,4]$. If one of $C_{0}$ or $C_{1}$ is |$\ket{0}$, then ancilla =|$\ket{0}$, and the target will not be flipped in Step III. If $C_{0}=C_{1}=\ket{1}$, but one of $C_{2},C_{3},C_{4}$ is |$\ket{0}$, then ancilla becomes |$\ket{1}$, and after applying X-gate on $C_{0},C_{1}$ will make, $C_{0}=C_{1}=\ket{0}$. If one of $C_{2},C_{3}$ is |$\ket{0}$, then $C_{0}$ and consequently, $C_{1}$ also remains at |$\ket{0}$ state. Similarly, if $C_{4}=\ket{0}$, then also $C_{1}$ remains |$\ket{0}$, and in both the scenario, the target will not be flipped in Step III.
In Section IV.2, we modify the design by introducing additional ancilla qubits in Step I and subsequently creating more conditionally clean ancilla to begin with, which eventually reduces the overall Toffoli depth of the complete circuit.
III Warm-up: A consolidated view on recent developments in Toffoli decomposition
As we have already explained, the decomposition of complex quantum gates into simpler and more practical quantum gate sets has been a topic of interest since the inception of quantum computing. More specifically, there have been substantial developments in the direction of multi-controlled Pauli (or, more precisely, MCT) decomposition in the last decades. This is a warm-up section where we provide a comprehensive overview of existing benchmarks for single and multi-controlled Toffoli gate decompositions. It emphasizes the state-of-the-art optimized results, to the best of our knowledge, in terms of T-Count, T-Depth, and ancilla requirements, as summarized in Tables 1 and 2.
III.1 Single Toffoli Decomposition
To immediately dive into the technical issues, as shown in (nc, , Ch. 4, Sec. 4.3), there exists a Clifford + T decomposition of a single Toffoli gate using $7$ T gates and $9$ Clifford gates (2 H, 1 S, 6 CNOT), with a T-Depth of $6$ (Figure 3(a)). Through a simple manipulation of gate ordering, the T-Depth can be reduced to $4$ without altering the gate count. In 2013, Amy et al. (amy13, , Section 6) proposed the Toffoli decomposition (Figure 3(b)) with a T-Depth of $4$ that used one fewer Clifford gate and reduced the overall circuit depth from $12$ to $8$. In the same paper, the authors applied a meet-in-the-middle algorithm to present a Toffoli decomposition (Figure 3(c)) with a T-Depth of $3$ and an overall depth of $9$. For an exact decomposition of a single Toffoli gate (without measurement-based feedback), this is the lowest T-Depth that one can achieve without using any ancillary qubit.
Using one ancilla qubit, the authors also proposed a Toffoli decomposition circuit (Figure 4(a)) using $7$ T gates and $12$ Clifford gates, achieving a T-Depth $2$. Later, in the same year, Selinger (selinger13, , Section 2) proposed a Toffoli gate decomposition circuit (Figure 4(b)) with the lowest possible T-Depth of $1$, using $4$ additional ancilla qubits and $18$ Clifford gates.
(a) T-Depth: 6 nc .
(b) T-Depth: 4 amy13 .
(c) T-Depth: 3 amy13 .
Figure 3: Single Toffoli decomposition without using any ancilla, utilizing $7$ T gates.
(a) T-Depth: 2, Ancilla: 1 amy13 .
(b) T-Depth: 1, Ancilla: 4 selinger13 .
Figure 4: Single Toffoli decomposition with ancilla, utilizing $7$ T gates.
In selinger13 , Selinger also proposed a doubly-controlled $-i$X gate, which differs from the Toffoli gate only by a controlled-$\text{S}^{\dagger}$ gate between the two control qubits. In the same year, Jones jones13 modified this circuit using a measurement-based uncomputation technique to implement the exact Toffoli gate, utilizing a single ancilla qubit with a T-Count of 4 and a T-Depth of 1. The schematic diagram of the circuit has been shown in Figure 5.
Figure 5: Single Toffoli decomposition due to Jones jones13 .
Later, in 2019, Soeken integrated the designs of Selinger selinger13 and Jones jones13 , proposing a modified circuit for single Toffoli decomposition utilizing measurement-based updates jaques19 , as illustrated in Figure 6. This circuit requires one ancilla qubit, uses four T-gates with a T-Depth of 1, and has an overall depth of 8.
Figure 6: Single Toffoli decomposition due to Soeken jaques19 .
In 2018, Gidney introduced the concept of logical-AND gidney18 using measurement-based uncomputation, which has a T-Count of 4 and a T-Depth of 2, without using any additional ancilla. When multiple logical-AND circuits are employed within the same quantum circuit, all the initial T-gates used for preparing the state |$\ket{T}$, where |$\ket{T}=\text{TH}\ket{0}$, can be executed simultaneously at the beginning of the circuit, with a T-Depth of 1. Consequently, the effective T-Depth of the logical-AND decomposition reduces to 1. This is marked with (*) in the last row of the Table 1. To clarify the notation, we have
[…]
The circuit diagram for the logical-AND is presented in Figure 7.
Figure 7: Single Toffoli decomposition using logical-AND due to Gidney gidney18 .
 Reference  Ancilla  T-Count  T-Depth #Clifford  Depth
 Amy et al. amy13 0 7 4 8 8
 Amy et al. amy13 0 7 3 9 9
 Amy et al. amy13 1 7 2 12 11
 Selinger selinger13 4 7 1 18 8
 Jaques et al. jaques19 1 4 1 11 8
 Gidney gidney18 * 0 4 $1+1$ 9 9
Table 1: Decompositions of a single Toffoli gate. The table specifies the required ancilla qubits, T-Count, T-Depth, Clifford counts, and the overall depth for different decompositions. The circuit with the least T-Count is highlighted in blue, the least T-Depths in teal, and the decompositions without ancilla in red.
Table 1 summarizes various resource requirements (T-Count, T-Depth, and Ancilla count) for the Clifford + T decomposition of a single Toffoli gate.
III.2 Multi-Controlled Toffoli Decomposition
From Section II, it is known that the multi-controlled Pauli gates (X, Y, Z) can be transformed into one another using a constant number of Clifford gates, such as the Hadamard or the Phase gates, as described in Figure 1. As the primary focus of this work is to minimize the implementation cost of multi-controlled Toffoli (MCT) gates in terms of Toffoli count, Toffoli depth, and the ancilla count, the inclusion of additional Clifford gates does not affect the resource estimation. Therefore, the resource estimation for any of the aforementioned multi-controlled Pauli gates can be directly translated to others without requiring modification.
In 2021, Gidney and Jones gidney21 presented a construction (Figure 8) of a $3$-controlled Z gate using 6 T-gates having a T-Depth of 6. Additionally, they proposed that their design can be used for the construction of an $n$-controlled Pauli gate, with a T-Count of $64n-$ 6, using $n-2$ logical-AND gates.
Figure 8: CCCZ circuit using 6 T gates gidney21 .
Proposition 1.
Following the CCCZ circuit of gidney21 , an $n$-MCT gate can be constructed using $n-3$ ($2$-controlled) Toffoli, and a single $3$-controlled Toffoli (CCCX) gate with $n-2$ ancilla qubits, resulting in a Toffoli depth of $\lceil\log_{2}\frac{n}{3}\rceil+1^{*}$, where $1^{*}$ represents the depth of the CCCX gate. Replacing the Toffoli gates with logical-AND (Figure 7) yields a complete T-Depth of $\lceil\log_{2}\frac{n}{3}\rceil+6$. Additionally, the total Clifford count of the circuit is $169n-$ 16.
Proof.
As specified in gidney21 , out of the $n-2$ AND gates, the final AND gate and Toffoli gate can be merged to implement a CCCX gate, making the Toffoli count $(n-3)+1^{*}$, i.e., n-3 many ($2$-controlled) Toffoli and a single $3$-controlled Toffoli gate. Consider a tree data structure with the root node having 3-child nodes, and each of them forms a complete binary tree with a total of $n$ leaf nodes. Each of these binary trees has a depth $\lceil\log_{2}\frac{n}{3}\rceil$, and the root node, along with its three child nodes, contributes a depth of 1. Consequently, the T-Depth becomes $\lceil\log_{2}\frac{n}{3}\rceil+6$, where the CCCX gate contributes to the T-Depth of $6$. Since the Clifford count of each logical-AND is 9, and that of the CCCX gate 11, the total Clifford count becomes $169n-$ 16. ∎
In Oct 2024, Nakanishi et al. naka24 proposed a modified CCCZ circuit (Figure 9), reducing the T-Depth to 2 by utilizing an additional ancilla qubit while keeping a Clifford count of 14.
Figure 9: CCCZ circuit having T-Depth 2 naka24 .
Consequently, the T-Depth of the $n$-MCT decomposition is reduced to $\lceil\log_{2}\frac{n}{3}\rceil+2$, with the corresponding Clifford count given by $159n-$ 15 (see the second row of Table 2).
In Feb 2024, Nie et al. nie24 first used the notion of conditionally clean ancilla qubits derived from an existing (clean or dirty) ancilla and proposed a novel circuit decomposition (Figure 10) for the $n$-controlled Pauli gates using $\mathcal{O}(n)$ Toffoli gates. The outer layer of their MCT decomposition follows from gidney15 , and the inner layer has been parallelized to obtain an overall Toffoli depth of $\mathcal{O}(\log_{2}n)$, compared to the $\mathcal{O}(\log_{2}n)$ Toffoli depth in gidney15 . Additionally, by improving the design of a quantum incrementer, they developed an MCT circuit with a Toffoli count of $\mathcal{O}(n)$, and a Toffoli depth of $\mathcal{O}(\log^{2}n)$ without requiring any additional ancilla qubits. In this design, although the MCT decomposition does not use any additional ancilla, implementing the quantum incrementer requires one ancilla. Since, our primary focus here is to reduce Toffoli depth while increasing the clean ancilla count, we ignore the zero-ancilla implementation here.
Figure 10: MCT decomposition using conditionally clean ancilla due to nie24 .
Proposition 2.
The MCT circuit decomposition proposed by nie24 using a single clean ancilla, has a minimum Toffoli count of $44n+$ 4, and an exact Toffoli depth of $20\log_{2}n$.
Proof.
Step I and Step V implement 4-MCT gates following gidney15 , each requiring $4(4-2)=8$ Toffoli gates, all applied sequentially, resulting in a Toffoli depth of 16. Similarly, the 3-MCT implemented in Step III has a Toffoli count of 4 and a Toffoli depth of 4. Consequently, Steps I, III, and V together require a total of 20 Toffoli gates, with an overall Toffoli depth of 20, which is referred to in the paper as a Toffoli depth of $\mathcal{O}(1)$.
Additionally, in Step II, an $(n-4)$-MCT is decomposed into two $\left(\frac{n}{2}-2\right)$-MCT, each requiring a minimum of $42\left(\frac{n}{2}-2\right)=n-$ 4 Toffoli gates. Thus, the Toffoli count for Step II is $82n-$ 8. Similarly, Step IV also has a Toffoli count of $82n-$ 8. Therefore, the total Toffoli count for the entire process is given by $42(2n-8)+20=4n+$ 4. Moreover, from nie24 , we have $D(n)=D(n/2)+\mathcal{O}(n)$, and we estimated $\mathcal{O}(1)=20$, therefore, the exact Toffoli depth of the MCT circuit is $20\log_{2}n$. ∎
In July 2024, Khattar and Gidney khattar24 further optimized the MCT circuit implementations by leveraging the conditionally clean ancilla qubits to reduce the Toffoli depth while restricting the ancilla count to $1$ or $2$. They proposed an $n$-MCT circuit utilizing a single clean ancilla (Figure 11(a)), achieving a Toffoli count of $32n-$ 3 and a T-Count of $128n-$ 12. Since none of the Toffoli gates are applied simultaneously, the resulting Toffoli depth is $32n-$ 3, while the T-Depth is $32n-$ 3, following the T-Depth $1$ Toffoli implementation by jaques19 , requiring one more reusable ancilla.
Furthermore, when the availability of clean ancilla increases to $2$ (Figure 11(b)), the Toffoli depth reduces to $\mathcal{O}(\log_{2}n)$ while maintaining the Toffoli count constant. Figure 11 illustrates the circuit diagrams of $10$-MCT using $1$ and $2$ clean ancilla, due to khattar24 . Additionally, if the clean ancilla is replaced with the dirty ancilla, the Toffoli count increases to $3216n-$ 32, and the Toffoli depth doubles under both scenarios. As we focus here on the conditionally clean ancilla derived solely from clean ancilla, we do not delve into an exact analysis of the Toffoli depth for the dirty ancilla circuit implementation.
(a) Ancilla: 1, Toffoli depth: 17.
(b) Ancilla: 2, Toffoli depth: 13.
Figure 11: $10$-MCT circuit decompositions, each of them using 17 Toffoli gates khattar24 .
In this section, we provided the exact enumerations of Toffoli count and Toffoli depth of MCT circuit decomposition across various state-of-the-art results. For the exact enumeration of the Toffoli depth (and T-Depth), following khattar24 , we propose a novel approach of viewing the MCT decomposition with conditionally clean ancilla, which is a direct extension of khattar24 , thereby presented separately in Proposition 3 of Section IV.1. Later, the approach has also been used to demonstrate the trade-off between Toffoli depth and the clean ancilla count.
The state-of-the-art optimized results corresponding to $n$-controlled Toffoli decompositions have been summarized in Table 2.
Reference  Ancilla Toffoli count Toffoli depth  T-Count  T-Depth Gidney et al. gidney21 $n-2$ $(n-3)+1^{*}$ $\lceil\log_{2}\frac{n}{3}\rceil+1^{*}$ $64n-$ 6 $\lceil\log_{2}\frac{n}{3}\rceil+6$ Nakanishi et al. naka24 $n-1$ $(n-3)+1^{*}$ $\lceil\log_{2}\frac{n}{3}\rceil+1^{*}$ $64n-$ 6 $\lceil\log_{2}\frac{n}{3}\rceil+2$ Nie et al. nie24 1 $\geq 4n+4$ $20\log_{2}n$ $1616n+$ 16 $20\log_{2}n$ Khattar et al. khattar24 1 $32n-$ 3 $32n-$ 3 $128n-$ 12 $32n-$ 3 Khattar et al. khattar24 2 $32n-$ 3 $\approx 4\log_{2}n$ $128n-$ 12 $\approx 4\log_{2}n$ Ours $m_{1}(\ll n)+2$ $32n-m_{1}-$ 3 $\begin{cases}2\lfloor\log_{2}m_{1}\rfloor+4k,\text{if }n\in\left[\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k-1}\right.\\ \hskip 113.81102pt\left.+k-1,m_{1}2^{k}+k-2\right],\\ 2\lfloor\log_{2}m_{1}\rfloor+4k+2,\text{if }n\in\left[m_{1}2^{k}+k-1,\right.\\ \hskip 71.13188pt\left.\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k}+k-1\right],k\geq 2.\end{cases}$ $128n-4m_{1}-$ 12 Same as Toffoli depth
Table 2: Decompositions of $n$-controlled Toffoli gates. The table specifies the required ancilla qubits, Toffoli count, Toffoli depth, T-Count, and T-Depth. The circuit with the least T-Count is highlighted in blue, the least T-Depths in teal, and the decompositions with the least ancilla in red.
IV Further Reduction of Toffoli Depth
In this section, we adopt the MCT decomposition technique using the conditionally clean ancilla, as introduced in nie24 ; khattar24 , and propose a trade-off between the Toffoli depth with the clean ancilla, showing improvement over the recent work of Khattar and Gidney khattar24 . The authors proposed the MCT circuit decompositions techniques, achieving a Toffoli depth of $\mathcal{O}(n)$ with a single clean ancilla and $\mathcal{O}(\log_{2}n)$ with two clean ancillae, where in both cases, the Toffoli count becomes $32n-$ 3. Here, we revisit the MCT decompositions by khattar24 and provide an exact enumeration of the Toffoli depth (consequently T-Depth) that was not explicitly presented in khattar24 .
IV.1 Exact Enumeration of Toffoli Depth proposed by Khattar and Gidney (2024) in a Different Lens
In this subsection, we enumerate the exact Toffoli depth of an $n$-MCT circuit following khattar24 . The $n$-MCT decomposition utilizing a single ancilla requires $32n-$ 3 Toffoli gates, all applied sequentially, resulting in a Toffoli depth of $32n-$ 3. Subsequently, by employing the measurement-based uncomputation technique for Toffoli decomposition, as illustrated in Figure 6, the T-Depth is again the same, i.e., $32n-$ 3.
The $n$-MCT decomposition circuit khattar24 using two ancilla qubits consists of five steps. Step I and Step V combined have a Toffoli depth of 1 (using logical-AND). In Step II, information from the remaining $n-2$ control qubits is accumulated into $k(\ll n)$ qubits using the conditionally cleaned ancilla technique. Next, in Step III, a $(k+1)$-MCT is implemented using the unused ancilla, following the single-ancilla technique, having a Toffoli depth of $12(k+1)-3=2k-$ 1. In Step IV, the control qubits are returned to their original state through uncomputation, thereby having the same Toffoli count and depth as in Step II.
For an exact enumeration of the Toffoli depth, we introduce a new way of viewing Step I and Step II of the MCT circuit decomposition, as follows.
$2\rightarrow$ 11 11
${}_{+1}^{2}\rightarrow$ ${}_{+1}^{1}\rightarrow$ 11 11
${}_{+1}^{4}\rightarrow$ ${}_{+1}^{2}\rightarrow$ ${}_{+1}^{1}\rightarrow$ 11 11
${}_{+1}^{8}\rightarrow$ ${}_{+1}^{4}\rightarrow$ ${}_{+1}^{2}\rightarrow$ ${}_{+1}^{1}\rightarrow$ 11 11
${}_{+1}^{12}\rightarrow$ ${}_{+1}^{6}\rightarrow$ ${}_{+1}^{3}\rightarrow$ $2\rightarrow$ 11 11
where the first row corresponds to Step I, and the remaining rows correspond to Step II. Each layer of horizontal arrows represents a Toffoli depth of 1, and a row, $a\rightarrow\lceil a/2\rceil\rightarrow\ldots\rightarrow 1$ implies that the information from $a$ many qubits has been accumulated into a single qubit, with a Toffoli depth of $\lceil\log_{2}a\rceil$. Consequently, the first element of each row sums to $n$. Additionally, the number of layers ($k$) determines the size of the MCT gate required for Step III.
In the above example, Step I and Step II of a $32$-MCT decomposition have been shown using the conditionally cleaned ancilla technique. The Step I has a Toffoli depth of $1$, and Step II has a Toffoli depth of $7$. Further, in Step III, a $5$-MCT needs to be applied to hit the target qubit, having a Toffoli depth $2(5)-3=7$. However, a part of the smaller MCT is applied simultaneously with Step II and Step IV, thereby having an effective Toffoli depth of $3$ or $5$. Clearly, the uncomputation (Step IV) requires an additional Toffoli depth of $8$, i.e., the Toffoli depth of the complete circuit is $19$.
Using this new approach the $5$-MCT circuit using conditionally clean ancilla technique, presented in Figure 2, can be seen as follows.
$2\rightarrow$ 11 11
${}_{+1}^{2}\rightarrow$ ${}_{+1}^{1}\rightarrow$ $1,$
where the Toffoli depth for Step I and Step II combined is 3, which can also be verified from the actual circuit from Figure 2. Moreover, a $2$-controlled Toffoli needs to implemented in Step III to modify the target.
Remark 1.
The motivation for viewing the MCT decomposition using this new representation because it allows more efficient and accurate estimation of the Toffoli depth, compared to the standard circuit representation. Additionally, our approach provides insight into the number of simultaneous operations and the size of the MCT gate required for implementation in Step III. In the subsequent section, we present our MCT circuit design with additional ancilla qubits using the same representation.
Further, note that introducing additional rows in Step II, without increasing the overall Toffoli depth will increase the size of the MCT gate required in Step III. Consequently, the overall Toffoli depth of the circuit remains unchanged.
In the direction of enumerating exact Toffoli depth of an $n$-MCT decomposition due to khattar24 , using 2 clean ancilla, we have the following lemma.
Lemma 1.
The exact Toffoli depth, $\delta$ for Step II or Step IV of an $n$-MCT circuit decomposition, as proposed by khattar24 , using two clean ancilla qubits, varies with $n$ as follows.
$\delta=\begin{cases}2k-3,\,\text{ for }n\in\left[3\cdot 2^{k-2}+k-1,\,2^{k}+k-2\right],\\ 2k-2,\,\text{ for }n\in\left[2^{k}+k-1,\,3\cdot 2^{k-1}+k-1\right]\end{cases}$
where $k\in\mathbb{N}$, with $k\geq 2$.
Lemma 2.
The number of control qubits, $\sigma$, in the MCT gate implemented in Step III of an $n$-MCT circuit decomposition, as proposed by khattar24 , using two clean ancilla qubits, varies with $n$ as follows.
$\sigma=\left\{\begin{array}[]{lr}k&\,\,\text{for }n\in\left[3\cdot 2^{k-2}+k-1,\,2^{k}+k-1\right],\\ k+1&\,\,\text{for }n\in\left(2^{k}+k-1,\,3\cdot 2^{k-1}+k-1\right].\end{array}\right.$
where $k\in\mathbb{N}$, with $k\geq 2$. The corresponding Toffoli depths are $32k-$ 3, and $12(k+1)-3=2k-$ 1, respectively. However, a significant part of the smaller MCT is applied simultaneously with Step II and Step IV, making the effective Toffoli depth of Step III, either $3$ or $5$.
From above lemma, we can now estimate the exact Toffoli depth of the $n$-MCT circuit decomposition using 2 ancilla qubits, as proposed in khattar24 .
Proposition 3.
Following khattar24 , the exact Toffoli depth of an $n$-MCT decomposition using $2$ clean ancilla is lower bounded by $3\log_{2}n$.
Proof.
From the above lemma, the exact Toffoli depth, $\delta$, of the complete $n$-MCT circuit decomposition is given by
$\delta=\left\{\begin{array}[]{lr}1+2(2k-3)+3=4k-2&\\ \hskip 51.21504pt\text{for }n\in\left[3\cdot 2^{k-2}+k-1,\,2^{k}+k-1\right],\\ 1+2(2k-2)+3=4k&\\ \hskip 51.21504pt\text{for }n\in\left(2^{k}+k-1,\,3\cdot 2^{k-1}+k-1\right]\end{array}\right.$
where $k\in\mathbb{N}$, with $k\geq 2$. Since, here we are interested in showing the lower bound, we consider $n$ assumes the highest value in the range, and still show that $3\log_{2}n$ is strictly less than the corresponding depths, as follows.
$3\log_{2}(2^{k}+k-1)<3\log_{2}(2^{k+1})=3(k+1),$
which is less or equal to the depth $24k-$ 2, for $k\geq 5$.
$3\log_{2}(3\cdot 2^{k-1}+k-1)<3\log_{2}(4\cdot 2^{k-1})=3(k+1),$
which is less or equal to the depth $4k$ for $k\geq 3$. ∎
Although the above proposition establishes that the exact Toffoli depth using two clean ancilla, as per khattar24 , is lower bounded by $3\log_{2}n$, this bound is not tight and is closer to $4\log_{2}n$. In the following section, we analyze the trade-off between Toffoli depth and clean ancilla, demonstrating that with additional ancilla, the exact Toffoli depth can be reduced to $2\log_{2}n$.
IV.2 Exact Trade-Off between Toffoli Depth and Clean Ancillae
In this subsection, we explore the ancilla-Toffoli depth trade-off using the concept of conditionally clean ancillae and demonstrate improvements in the Toffoli depth compared to khattar24 by introducing additional ancilla qubits into the circuit while keeping the Toffoli count constant. Furthermore, for an $n$-controlled Toffoli gate with $m$ ancilla qubits, we propose an algorithm that determines the Toffoli depth (and consequently the T-Depth) for the MCT circuit decomposition and presents a graph illustrating the trade-off.
Construction 1.
Given $m$ ancilla qubits to implement an $n$-controlled Toffoli gate, we distribute $m$ as $m_{1}+m_{2}$, where $m_{1}$ ancilla qubits are allocated for Step I, and $m_{2}$ ancilla qubits are reserved for Step III. For $m=3$, we set $m_{1}=2$ and $m_{2}=1$. Furthermore, for $m\geq 4$, we assume $m_{2}=2$ to implement the smaller MCT in Step III using the $2$-clean ancilla technique described in khattar24 .
The circuit design is as follows.
  • In Step I, the $m_{1}$ Toffoli gates are applied to collect information from $2m_{1}$ control qubits and store them in $m_{1}$ ancilla qubits. We implement them using the logical-AND circuit to avoid resource requirements for uncomputation in the final step.
  • In Step II, information from the remaining control qubits is accumulated into $k(<n)$ qubits using the conditionally cleaned ancilla.
  • In Step III, a $(k+m_{1})$-controlled Toffoli gate is implemented using the remaining $m_{2}$ many ancilla, following the $2$-ancilla technique, and the target qubit is modified.
  • In Step IV, uncomputation is performed to return the control qubits to their original state. As earlier, the Toffoli count and Toffoli depth of Step IV are identical to those in Step II.
  • Finally, in Step V, the first $m_{1}$ many Toffoli gates are uncomputed to clean up the ancilla, using logical-AND based uncomputation. This step reduces the overall Toffoli count by $m_{1}$ compared to the circuit presented in khattar24 .
The schematic diagram of the MCT decomposition using $m=m_{1}+m_{2}$ many ancilla is similar to the diagram we used for the exact Toffoli depth enumeration in Section IV.1.
$2m_{1}\rightarrow$ $m_{1}$
[…]
Lemma 3.
For the proof of correctness, the circuit decomposition presented in Construction 1 is an $n$-MCT.
Proof.
In the above MCT decomposition, if any of the first $2m_{1}$ qubits is in the |$\ket{0}$ state, one of the $m_{1}$ ancilla qubits will also remain in |$\ket{0}$, preventing the target flip in Step III. Similarly, if any of the $n-2m_{1}$ qubits is |$\ket{0}$, the corresponding row’s end qubit remains in |$\ket{0}$, ensuring the target is not flipped in Step III. This proves that the designed circuit correctly implements an $n$-MCT gate. ∎
Example 1.
For $m_{1}=3$, the $32$-controlled Toffoli can be viewed as:
$6\rightarrow$ 33 33
${}_{+1}^{6}\rightarrow$ ${}_{+1}^{3}\rightarrow$ $2\rightarrow$ 11 11
$12\rightarrow$ $6\rightarrow$ $3\rightarrow$ ${}_{+1}^{1}\rightarrow$ 11 11
${}_{+1}^{6}\rightarrow$ ${}_{+1}^{3}\rightarrow$ $2\rightarrow$ 11 11
where the Toffoli depth for Step I is $1$, and the Toffoli depth for Step II (and Step IV) are $5$ each. Additionally, in Step III, we need to implement a $6$-controlled Toffoli using $m_{2}=2$ ancilla, which requires an additional Toffoli depth of $3$. Consequently, the Clifford + Toffoli decomposition of the $32$-controlled Toffoli gate, utilizing $(3+2)=5$ ancilla qubits, achieves a total Toffoli depth of $1+2(5)+3=14$. In contrast, the $32$-controlled Toffoli decomposition with $2$ ancilla qubits, as presented in khattar24 , results in a Toffoli depth of $19$.
Similarly, for $m_{1}=4$, the $32$-controlled Toffoli can be viewed as:
$8\rightarrow$ 44 44
${}_{+1}^{8}\rightarrow$ ${}_{+1}^{4}\rightarrow$ ${}_{+1}^{2}\rightarrow$ ${}_{+1}^{1}\rightarrow$ 11 11
${}_{+1}^{14}\rightarrow$ ${}_{+1}^{7}\rightarrow$ $4\rightarrow$ $2\rightarrow$ 11 11
where the Toffoli depth remains $14$. Thus, it can be observed that beyond a certain point, further increases in the number of ancilla qubits do not necessarily lead to a reduction in the Toffoli depth when using conditionally clean ancilla.
The Toffoli depth of the MCT decomposition using $m=m_{1}+m_{2}$ many ancilla, can be estimated from the following lemma.
Lemma 4.
The exact Toffoli depth of the Step II or Step IV of an $n$-MCT circuit decomposition, implemented above, using $m=m_{1}+m_{2}$ clean ancilla following the conditionally clean ancilla technique, varies with $n$ as follows.
$\delta=\begin{cases}\lfloor\log_{2}m_{1}\rfloor+2k-3,\\ \indent\text{for }n\in\left[\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k-1}+k-1,\right.\\ \hskip 147.95424pt\left.m_{1}2^{k}+k-2\right]\\ \lfloor\log_{2}m_{1}\rfloor+2k-2,\\ \indent\text{for }n\in\left[m_{1}2^{k}+k-1,\right.\\ \hskip 73.97733pt\left.\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k}+k-1\right]\end{cases}$
where $k\in\mathbb{N}$, with $k\geq 2$.
Lemma 5.
The number of control qubits, $\sigma$, in the Step III of an $n$-MCT circuit decomposition, implemented above using $m=m_{1}+m_{2}$ ancilla qubit, varies with $n$ as follows.
$\sigma=\left\{\begin{array}[]{l}k+m_{1}-1,\\ \indent\text{ for }n\in\left[\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k-1}+k-1,\right.\\ \hskip 147.95424pt\left.m_{1}2^{k}+k-1\right]\\ k+m_{1},\\ \indent\text{ for }n\in\left(m_{1}2^{k}+k-1,\right.\\ \hskip 73.97733pt\left.\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k}+k-1\right]\end{array}\right.$
where $k\in\mathbb{N}$, with $k\geq 2$. The corresponding Toffoli depths are $4\log_{2}(k+m_{1}-1)$, and $4\log_{2}(k+m_{1})$, respectively. However, a significant part of the smaller MCT is applied simultaneously with Step II and Step IV, making the effective Toffoli depth of Step III, either $3$ or $5$.
From above lemma, we can now estimate the exact Toffoli depth of the $n$-MCT circuit decomposition using $m=m_{1}+m_{2}$ ancilla qubits, as presented above.
Theorem 1.
The exact Toffoli depth, $\delta$, of the complete $n$-MCT circuit decomposition using $m=m_{1}+m_{2}$ clean ancilla is given by
$\delta=\begin{cases}2\lfloor\log_{2}m_{1}\rfloor+4k,\\ \indent\text{for }n\in\left[\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k-1}+k-1,\right.\\ \hskip 147.95424pt\left.m_{1}2^{k}+k-2\right]\\ 2\lfloor\log_{2}m_{1}\rfloor+4k+2,\\ \indent\text{for }n\in\left[m_{1}2^{k}+k-1,\right.\\ \hskip 73.97733pt\left.\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k}+k-1\right]\end{cases}$
where $k\in\mathbb{N}$, with $k\geq 2$.
Proof.
From Lemma 4, and 5, the exact Toffoli depth of an $n$-MCT, from Step I-IV, for $n\in\left[\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k-1}+k-1,m_{1}2^{k}+k-2\right]$ is
$1+2\left(\lfloor\log_{2}m_{1}\rfloor+2k-3\right)+5=2\lfloor\log_{2}m_{1}\rfloor+4k.$
For $n\in\left[m_{1}2^{k}+k-1,\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k}+k-1\right]$, the exact Toffoli depth of the $n$-MCT becomes
$21+2\left(\lfloor\log_{2}m_{1}\rfloor+2k-2\right)+5=2\lfloor\log_{2}m_{1}\rfloor+4k+$ 2 .
It is now understood that the depth of $n$-MCT can be achieved in $\mathcal{O}(\log_{2}n)$, which can be explicitly written $c\log_{2}n$, where $c$ needs to be properly estimated, for exact trade-offs which is the main motivation in this paper. In Proposition 3, we have shown that in the MCT decomposition using two clean ancilla with the technique from khattar24 , $c$ is strictly greater than 3, in fact, it is around 4. However, using the ancilla - Toffoli depth trade-off, $c$ can be further reduced to a certain extent. In the following subsection, we demonstrate that with the conditionally clean ancilla technique, $c$ always remains strictly greater than 1, regardless of the number of ancilla used.
IV.3 Proving the Lower Bound on Toffoli Depth using Conditionally Clean Ancilla
In this subsection, we show that the Toffoli depth of an $n$-MCT using the conditionally clean ancilla technique can never be reduced to $\lceil\log_{2}n\rceil$.
Theorem 2.
This is in reference to Construction 1 for $n$-MCT.
  1. 1.
    Assuming that the control states do not need to be restored to their original values, using the conditionally clean ancilla technique, the exact Toffoli depth must be strictly greater than $\lceil\log_{2}n\rceil$, irrespective of the number of available ancilla.
  2. 2.
    When the control qubits are required to be returned to their original state upon completion, the Toffoli depth becomes strictly greater than $2\lceil\log_{2}n\rceil$.
Proof.
From Proposition 1, it is evident that the Toffoli depth remains constant over a range of values for $n$. Since, here we are interested in showing the lower bound, we consider $n$ assumes the highest value in the range, and still show that $\log_{2}n$ is strictly less than the corresponding depth, as follows.
Let us first proof the item 1. As the control states are not required to be returned to their original state, we consider the Toffoli depth due to Step I, Step II, and half of Step III only, which is
$\delta^{\prime}=\begin{cases}\lfloor\log_{2}m_{1}\rfloor+2k+1,\text{if }n\in\left[\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k-1}\right.\\ \hskip 113.81102pt\left.+k-1,m_{1}2^{k}+k-2\right]\\ \lfloor\log_{2}m_{1}\rfloor+2k+2,\text{if }n\in\left[m_{1}2^{k}+k-1,\right.\\ \hskip 85.35826pt\left.\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k}+k-1\right]\end{cases}$
Thus, we need to show the following two cases:
  • Case I: $\lceil\log_{2}\left(m_{1}2^{k}+k-2\right)\rceil\leq\lfloor\log_{2}m_{1}\rfloor+2k+1$,
  • Case II: $\lceil\log_{2}\left(\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k}+k-1\right)\rceil\leq\lfloor\log_{2}m_{1}\rfloor+2k+2$.
Case I: Showing $\lceil\log_{2}\left(m_{1}2^{k}+k-2\right)\rceil\leq\lfloor\log_{2}m_{1}\rfloor+2k+1$ is equivalent to showing $m_{1}2^{k}+k-2\leq 2^{\lfloor\log_{2}m_{1}\rfloor+2k}$. We prove this by induction on $k\geq 2$, for $k\in\mathbb{N}$.
Base case: For $k=2$, RHS = $2^{\lfloor\log_{2}m_{1}\rfloor+4}>2^{\log_{2}m_{1}-1+4}=8m_{1}$, which is strictly greater than $4m_{1}$, LHS for $k=2$.
Induction hypothesis: Suppose the result holds for $k=k_{1}\in\mathbb{N}$, i.e.,
$m_{1}2^{k_{1}}+k_{1}-2\leq 2^{\lfloor\log_{2}m_{1}\rfloor+2k_{1}}.$
Inductive step: We need to show that the result holds for $k=k_{1}+1$, i.e.,
$m_{1}2^{k_{1}+1}+k_{1}-1\leq 2^{\lfloor\log_{2}m_{1}\rfloor+2(k_{1}+1)}.$
LHS, $m_{1}2^{k_{1}+1}+k_{1}-1=2\left(m_{1}2^{k_{1}}+k_{1}-2\right)-k_{1}+3\leq 2\left(2^{\lfloor\log_{2}m_{1}\rfloor+2k_{1}}\right)-k_{1}+3,$ which is strictly less than $4\left(2^{\lfloor\log_{2}m_{1}\rfloor+2k_{1}}\right)=2^{\lfloor\log_{2}m_{1}\rfloor+2(k_{1}+1)}$, RHS.
Therefore, Case I holds for all $k\geq 2$ for some $k\in\mathbb{N}$.
Case II: Showing $\lceil\log_{2}\left(\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k}+k-1\right)\rceil\leq\lfloor\log_{2}m_{1}\rfloor+2k+2$ is equivalent to showing
$\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k}+k-1\leq 2^{\lfloor\log_{2}m_{1}\rfloor+2k+1}.$
We again prove this by induction on $k\geq 2$, for $k\in\mathbb{N}$.
Base case: For $k=2$, LHS = $4\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)+1=4m_{1}+1+2^{\lfloor\log_{2}m_{1}\rfloor+1}$. Similarly, for $k=2$, RHS = $2^{\lfloor\log_{2}m_{1}\rfloor+4+1}=16\cdot 2^{\lfloor\log_{2}m_{1}\rfloor+1}$, which can be distribute into $15\cdot 2^{\lfloor\log_{2}m_{1}\rfloor+1}+\cdot 2^{\lfloor\log_{2}m_{1}\rfloor+1}$.
Now, $115\cdot 2^{\lfloor\log_{2}m_{1}\rfloor+1}>15\cdot 2^{\log_{2}m_{1}-1+1}=15m_{1}>4m_{1}+$ 1. Thus, the base case holds.
Induction hypothesis: Suppose the result holds for $k=k_{1}\in\mathbb{N}$, i.e.,
$\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k_{1}}+k_{1}-1\leq 2^{\lfloor\log_{2}m_{1}\rfloor+2k_{1}+1}.$
Inductive step: We need to show that the result holds for $k=k_{1}+1$, i.e.,
$\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k_{1}+1}+k_{1}\leq 2^{\lfloor\log_{2}m_{1}\rfloor+2k_{1}+3}.$
LHS = $\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k_{1}+1}+k_{1}$
$=$ $2\displaystyle 2\left[\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k_{1}}+k_{1}-1\right]-k_{1}+$ 2
$\leq$ 2($2\displaystyle 2\left(2^{\lfloor\log_{2}m_{1}\rfloor+2k_{1}+1}\right)-k_{1}+$ 2
$<$ $4\left(2^{\lfloor\log_{2}m_{1}\rfloor+2k_{1}+1}\right)=2^{\lfloor\log_{2}m_{1}\rfloor+2k_{1}+3}=\text{ RHS}.$
Therefore, Case II holds for all $k\geq 2$ for some $k\in\mathbb{N}$. In conclusion, using the conditionally clean ancilla technique, the exact Toffoli depth of an $n$-controlled Toffoli decomposition can not be reduced to $\lceil\log_{2}n\rceil$, even when the control qubits are not restored to their original states.
Item 2, where the control qubits are restored to their original state, can be proved inductively in a similar manner. The factor of 2 needs to be multiplied in this case because the full depth, $\delta$, from Proposition 3, is related to $\delta^{\prime}$ as $\delta=2\delta^{\prime}-2$, i.e., almost twice the depth considered in the first part. ∎
The above theorem shows that using the conditionally clean ancilla, the constant factor $c$ can never be reduced to $1$. In the next section, we demonstrate that the exact Toffoli depth in the Clifford + Toffoli decomposition of an $n$-MCT, is lower bounded by $\lceil\log_{2}n\rceil$, regardless of the technique or the number of ancilla qubits used. This bound is achievable through complete binary tree decomposition of $n$-controlled Toffoli gates.
V Tight $\lceil\log_{2}n\rceil$ Lower Bound on Toffoli Depth
In this section, we show in a more general framework that the exact Toffoli depth in the Clifford + Toffoli decomposition of an $n$-controlled Toffoli gate is lower bounded by $\lceil\log_{2}n\rceil$, which is exactly $\log_{2}n$ when $n=2^{k}$ for some $k\in\mathbb{N}$. Additionally, the exact T-Depth also becomes $\lceil\log_{2}n\rceil$, utilizing $22n-$ 2 ancilla qubits and a T-Count of $4(n-1)$, provided by jones13 . Alternatively, following Gidney’s logical-AND circuit gidney18 , the ancilla count can be reduced to $n-2$, maintaining the same T-Count, while the T-Depth increases to $\lceil\log_{2}n\rceil+1$.
Theorem 3.
The exact Toffoli depth in the Clifford + Toffoli decomposition of an $n$-MCT is lower bounded by $\lceil\log_{2}n\rceil$, given any technique or any number of ancilla qubits used.
Proof.
Each $2$-controlled Toffoli gate encodes the information of two control qubits into one target qubit. This process can be visualized as a binary tree, where the leaf nodes represent the $n$ control qubits, and their information is successively accumulated in internal parent nodes.
To implement an $n$-MCT, we begin with $n$ leaf nodes, where each pair of control qubits is combined using a Toffoli gate, reducing the number of active qubits in each round. At each stage, either two parent nodes or one parent node along with a leftover node from the previous round are further combined into a new parent node. This process continues iteratively until the final Toffoli gate transfers the accumulated information to the root node (the target qubit).
Since each Toffoli gate reduces two qubits into one, the number of required levels in the binary tree is given by the height of a binary tree with $n$ leaf nodes. The depth of such a binary tree is at least $\lceil\log_{2}n\rceil$. Hence, the Toffoli depth of the $n$-MCT is also at least $\lceil\log_{2}n\rceil$, proving the claim. ∎
A complete binary tree structure, having $n$ leaf nodes, has a hight $\log_{2}n$, i.e., when $n$ is not a power of $2$. When $n=2^{k}$, for some $k\in\mathbb{N}$, then it becomes a perfect binary tree, having the depth exactly $\log_{2}n=k$. In both the cases, the number of internal nodes, i.e., ancilla qubit is $n-2$. Additionally, the Toffoli count for both these cases becomes $n-1$.
For example, a $32$-MCT can be implemented using $30$ ancilla qubits, with $33$ Toffoli gates, having a minimum Toffoli depth of $5$. Similarly, all $n$-MCT decomposition, with $17\leq n\leq 32$, can be implemented with a minimum Toffoli depth of $5$. Figure 12 provides a schematic diagram of the $7$-MCT circuit decomposition using $5$ ancilla qubits, and $6$ Toffoli gates, with a Toffoli depth of $\lceil\log_{2}7\rceil=3$. In the diagram, the first four Toffoli gates in the first two columns are implemented simultaneously, having depth 1. The next two Toffoli gates are applied sequentially, having Toffoli depth 1, each.
Figure 12: $7$-MCT circuit decomposition with a Toffoli depth 3.
In this context, we present the following corollaries concerning the lower bound on the T-depth of an $n$-MCT circuit implementation via Clifford+Toffoli decomposition, regardless of the number of ancilla qubits used. These results can be obtained by first constructing an $\lceil\log_{2}n\rceil$ Toffoli depth circuit for the $n$-MCT decomposition, followed by further decomposing the Toffoli gates into Clifford+T gates as outlined in Table 1.
Corollary 1.
Using the measurement-based Toffoli decomposition circuit proposed by jaques19 , the T-depth of the Clifford+T decomposition of an $n$-MCT gate, implemented via Clifford+Toffoli decomposition, is lower bounded by $\lceil\log_{2}n\rceil$. Furthermore, the circuit requires $22n-$ 2 ancilla qubits, and the T-count is $44n-$ 4.
Corollary 2.
Using the logical-AND circuit proposed by gidney18 , an $n$-MCT gate can be implemented with a Clifford+T decomposition utilizing $n-2$ ancilla qubits. The resulting circuit has a T-depth of $\lceil\log_{2}n\rceil+1$ and a T-count of $44n-$ 4.
Towards, the conclusion, let us outline a generalized approach that may provide a theoretical understanding. Given the Algebraic Normal Form (ANF) of a Boolean function, $f:\{0,1\}^{n}\rightarrow\{0,1\}^{m}$, naturally the maximum algebraid degree can be $n$. Thus, the reversible quantum circuit implementing $f$, can be designed with a Toffoli depth of $\lceil\log_{2}n\rceil$, utilizing an exponential number of ancilla qubits and $\mathcal{O}(n)$ Clifford gates. This is because, in the ANF, the maximum degree is $n$, which means that $n$ different inputs are ANDed. Thus, an $n$-MCT is enough. Further, the ANF may contain at most $2^{n}$ terms, and that may require exponential number of ancilla qubits. This is achieved by implementing all the required multi-controlled Toffoli (MCT) gates in parallel to compute the non-linear terms of the ANF. This follows from Theorem 3.
This shows that any $n$-degree Boolean function can be implemented with a Toffoli depth of $\lceil\log_{2}n\rceil$. This idea may also be used while considering the cryptanalytic techniques, when the quantum circuits are actually implemented. In this direction, it is understood that the depth may really be quite less for each module to be implemented. Regarding the anicillae, for practical purposes, the ANF contains poly$(n)$ many terms in it, and in such cases the number of ancillae will also be poly$(n)$, instead of the generic exponential bound.
VI Conclusion
In this paper, we revisited the $n$-controlled Toffoli decomposition using the conditionally clean ancilla technique described by Khattar and Gidney khattar24 and proposed an exact trade-off between Toffoli depth and the availability of clean ancilla qubits. By leveraging additional ancilla qubits, we achieved a lower Toffoli depth compared to their approach. Furthermore, we demonstrated that the conditionally clean ancilla technique cannot reduce the Toffoli depth strictly to $\lceil\log_{2}n\rceil$, irrespective of unlimited availability of clean ancilla.
Additionally, we established that, regardless of the decomposition technique or available ancillae, the Toffoli depth of an $n$-MCT circuit is fundamentally lower-bounded by $\lceil\log_{2}n\rceil$, with the optimal depth achieved through binary tree-based MCT decomposition. Finally, by incorporating Soeken’s measurement-based uncomputation technique (as referred in jaques19 ) for Toffoli decomposition, we extended this lower bound to T-depth as well.

### v2

@@
 Unlike classical gates, quantum gates are inherently reversible and are mathematically represented by unitary matrices.
-The doubly-controlled X-gate, commonly known as the Toffoli gate, is among the most significant quantum gates, with critical applications in arithmetic operations, reversible computing, and oracle constructions.
+The doubly-controlled X-gate, commonly known as the Toffoli gate, is among the most significant quantum gates, with critical applications in arithmetic operations, reversible computing, and oracle constructions wang2024boosting ; wang2024optimal ; wang2025comprehensive ; wang2024Minimum ; wang2024Reducing ; wang2023higher .
 However, the Toffoli gate’s high resource demands, particularly in terms of T-Count, T-Depth, and ancilla qubits, can greatly influence the efficiency of fault-tolerant quantum circuits.
@@
 An $n$-controlled Toffoli is an $(n+1)$-qubit gate, where $n$ many qubits control the outcome of a single target qubit.
-Like ($2$-controlled) Toffoli, Multi Controlled Toffoli (MCT) also has a huge applications in quantum arithmetic, error correction, and implementation of quantum algorithms.
+Like ($2$-controlled) Toffoli, Multi Controlled Toffoli (MCT) also has a huge application in quantum arithmetic, error correction, and implementation of quantum algorithms.
 For example, in Grover’s Search, implementing the $n$-bit AND function requires an $n$-controlled Toffoli ($n$-MCT) gate.
@@
 It is well known that decomposing/ designing larger circuits with smaller components requires additional qubits, known as ancilla qubits, to reduce depth.
-The important task for optimizing the number of gates, the depth and the number of additional qubits has received serious attention for a long time and one may refer to the seminal work of Moore and Nilsson moore01 two decades back.
-A more recent and comprehensive discussion is available in jiang20 and we like to quote from that work “Can we characterize the relationship between the number of ancilla and the possible optimal depth?” The paper jiang20 is concerned with CNOT, whereas in this paper we concentrate on Toffoli.
+The important task of optimizing the number of gates, the depth, and the number of additional qubits has received serious attention for a long time, and one may refer to the seminal work of Moore and Nilsson moore01 two decades back.
+A more recent and comprehensive discussion is available in jiang20 , and we like to quote from that work, “Can we characterize the relationship between the number of ancilla and the possible optimal depth?” The paper jiang20 is concerned with CNOT, whereas in this paper we concentrate on Toffoli.
 However, the fundamental motivation of quantum circuit synthesis revolves around this very question.
@@
 More recent works nie24 ; khattar24 have introduced the conditionally clean ancilla technique, which significantly reduces both the Toffoli depth and ancilla count in the MCT decomposition.
-Before proceeding further let us explain the concept of conditionally clean ancilla that will be repeatedly referred in this work.
+Before proceeding further, let us explain the concept of conditionally clean ancilla that will be repeatedly referred to in this work.
 Generally, ancilla qubits are of two types, namely the clean ancilla, which is initialized to |$\ket{0}$ at the beginning of the circuit, and the dirty ancilla, which is initialized to some unknown quantum state at the beginning.
 The standard practice is to re-initialize the ancilla qubit to its initial state at the end of the computation.
-Conditionally clean ancilla nie24 ; khattar24 are the set of working (control) qubits, derived form the existing ones, that are treated as the clean ancilla based on certain assumptions, and re-initialized at the end of the computation.
-Although, earlier works nie24 ; khattar24 did consider the dirty ancilla, in this paper, we are focusing on the conditionally clean ancilla derived from the clean ones only.
+Conditionally clean ancilla nie24 ; khattar24 is the set of working (control) qubits, derived from the existing ones, that are treated as the clean ancilla based on certain assumptions, and re-initialized at the end of the computation.
+Although earlier works nie24 ; khattar24 did consider the dirty ancilla, in this paper, we are focusing on the conditionally clean ancilla derived from the clean ones only.
 Very recently, the conditionally clean ancilla technique has gained significant attention, including its applications in quantum adders remaud25 .
-The resource optimization metrics under consideration are the Toffoli count , Toffoli depth and Ancilla count.
+The resource optimization metrics under consideration are the Toffoli count, Toffoli depth, and Ancilla count.
 The Toffoli count and Toffoli depth are further refined with the T-Count and T-Depth as one may refer to Table 1.
@@
 Additionally, we also explain the recent developments of MCT decompositions describing the existing best results in terms of Toffoli count, Toffoli depth, and ancilla, which can be subsequently decomposed into T-Count and T-Depth, as summarized in Table 2.
-Given that there are several developments in very recent times, this section provide a holistic view of the existing results.
+Given that there are several developments in very recent times, this section provides a holistic view of the existing results.
 Based on this, we present our contributions.
 In section IV, we explore the trade-off between the (clean) ancilla and the Toffoli depth using the existing techniques related to conditionally clean ones.
-In Section IV.1, we take a different look in viewing the MCT circuit decomposition using conditionally clean ancilla technique of khattar24 , and enumerate their exact Toffoli count.
+In Section IV.1, we take a different look at viewing the MCT circuit decomposition using the conditionally clean ancilla technique of khattar24 and enumerate their exact Toffoli count.
 Following the trade-off, we show the reduction in Toffoli depth, and therefore in T-Depth (Construction 1, Example 1) compared to the recent work of Khattar and Gidney khattar24 , by introducing additional clean ancilla into the circuit, while keeping the Toffoli count constant.
@@
 Moreover, using the logical-AND circuit by Gidney gidney18 , the ancilla count can be reduced to $n-2$ keeping the T-Count constant.
-However, the exact T-Depth in case of gidney18 becomes one more, i.e., $\lceil\log_{2}n\rceil+1$.
+However, the exact T-Depth in the case of gidney18 becomes one more, i.e., $\lceil\log_{2}n\rceil+1$.
 To highlight our contribution, we are looking at the exact counts and depth instead of their complexity order.
@@
 […]
-However, the Clifford gates alone do not constitute a universal set of quantum gates because certain gates.
-For example, the T $=\sqrt{\text{S}}$ gate, cannot be arbitrarily approximated using only Clifford operations.
+However, the Clifford gates alone do not constitute a universal set of quantum gates because certain gates, for example, the T $=\sqrt{\text{S}}$ gate, cannot be arbitrarily approximated using only Clifford operations.
 Therefore, the Clifford group, when augmented with the T gate, forms a universal quantum gate set for quantum computation.
@@
 In quantum computing, the Multi-Controlled Toffoli (MCT) gates play a crucial role in the design of complex quantum algorithms, including error correction codes and arithmetic operations.
-Despite its utility, the implementation of MCT gates exploiting the Clifford + T gate set, requires careful decomposition to optimize resource usage, such as minimizing the T-Count, T-Depth, and the number of ancillary qubits.
+Despite its utility, the implementation of MCT gates exploiting the Clifford + T gate set requires careful decomposition to optimize resource usage, such as minimizing the T-Count, T-Depth, and the number of ancillary qubits.
 These optimizations are essential for practical quantum computation, where resource efficiency is a critical consideration.
@@
 If the clean ancilla is reversed, it means both the control qubits were |$\ket{1}$; thus, applying X gates on the control qubits will change them to |$\ket{0}$ and consequently can be used as conditionally clean ancilla in the next rounds.
-Similarly, in the following round, additional Toffoli gates are implemented targeting these conditionally clean ancilla and applying the X gate on the control qubits will again make them conditionally clean for the next round, and the process continues until we exhaust all the control qubits.
+Similarly, in the following round, additional Toffoli gates are implemented targeting these conditionally clean ancilla, and applying the X gate on the control qubits will again make them conditionally clean for the next round, and the process continues until we exhaust all the control qubits.
 Once the information from all the control qubits is accumulated in a few qubits, another (smaller) MCT gate is implemented with these qubits along with the ancilla as the control qubits and the original target as its target, so that if the ancilla was not modified in the first step, the target will also not be modified.
@@
 If one of $C_{2},C_{3}$ is |$\ket{0}$, then $C_{0}$ and consequently, $C_{1}$ also remains at |$\ket{0}$ state.
-Similarly, if $C_{4}=\ket{0}$, then also $C_{1}$ remains |$\ket{0}$, and in both the scenario, the target will not be flipped in Step III.
+Similarly, if $C_{4}=\ket{0}$, then also $C_{1}$ remains |$\ket{0}$, and in both scenario, the target will not be flipped in Step III.
 In Section IV.2, we modify the design by introducing additional ancilla qubits in Step I and subsequently creating more conditionally clean ancilla to begin with, which eventually reduces the overall Toffoli depth of the complete circuit.
@@
 In this design, although the MCT decomposition does not use any additional ancilla, implementing the quantum incrementer requires one ancilla.
-Since, our primary focus here is to reduce Toffoli depth while increasing the clean ancilla count, we ignore the zero-ancilla implementation here.
+Since our primary focus here is to reduce Toffoli depth while increasing the clean ancilla count, we ignore the zero-ancilla implementation here.
 Figure 10: MCT decomposition using conditionally clean ancilla due to nie24 .
 Proposition 2.
-The MCT circuit decomposition proposed by nie24 using a single clean ancilla, has a minimum Toffoli count of $44n+$ 4, and an exact Toffoli depth of $20\log_{2}n$.
+The MCT circuit decomposition proposed by nie24 using a single clean ancilla has a minimum Toffoli count of $44n+$ 4 and an exact Toffoli depth of $20\log_{2}n$.
 Proof.
@@
 Clearly, the uncomputation (Step IV) requires an additional Toffoli depth of $8$, i.e., the Toffoli depth of the complete circuit is $19$.
-Using this new approach the $5$-MCT circuit using conditionally clean ancilla technique, presented in Figure 2, can be seen as follows.
+Using this new approach, the $5$-MCT circuit using a conditionally clean ancilla technique, presented in Figure 2, can be seen as follows.
 $2\rightarrow$ 11 11
@@
 where the Toffoli depth for Step I and Step II combined is 3, which can also be verified from the actual circuit from Figure 2.
-Moreover, a $2$-controlled Toffoli needs to implemented in Step III to modify the target.
+Moreover, a $2$-controlled Toffoli needs to be implemented in Step III to modify the target.
 Remark 1.
@@
 Consequently, the overall Toffoli depth of the circuit remains unchanged.
-In the direction of enumerating exact Toffoli depth of an $n$-MCT decomposition due to khattar24 , using 2 clean ancilla, we have the following lemma.
+In the direction of enumerating the exact Toffoli depth of an $n$-MCT decomposition due to khattar24 , using 2 clean ancilla, we have the following lemma.
 Lemma 1.
@@
 However, a significant part of the smaller MCT is applied simultaneously with Step II and Step IV, making the effective Toffoli depth of Step III, either $3$ or $5$.
-From above lemma, we can now estimate the exact Toffoli depth of the $n$-MCT circuit decomposition using 2 ancilla qubits, as proposed in khattar24 .
+From the above lemma, we can now estimate the exact Toffoli depth of the $n$-MCT circuit decomposition using 2 ancilla qubits, as proposed in khattar24 .
 Proposition 3.
@@
 where $k\in\mathbb{N}$, with $k\geq 2$.
-Since, here we are interested in showing the lower bound, we consider $n$ assumes the highest value in the range, and still show that $3\log_{2}n$ is strictly less than the corresponding depths, as follows.
+Since here, we are interested in showing the lower bound, we consider $n$ assumes the highest value in the range and still show that $3\log_{2}n$ is strictly less than the corresponding depths, as follows.
 $3\log_{2}(2^{k}+k-1)<3\log_{2}(2^{k+1})=3(k+1),$
@@
 which is less or equal to the depth $4k$ for $k\geq 3$. ∎
-Although the above proposition establishes that the exact Toffoli depth using two clean ancilla, as per khattar24 , is lower bounded by $3\log_{2}n$, this bound is not tight and is closer to $4\log_{2}n$.
+Although the above proposition establishes that the exact Toffoli depth using two clean ancillae, as per khattar24 , is lower bounded by $3\log_{2}n$, this bound is not tight and is closer to $4\log_{2}n$.
 In the following section, we analyze the trade-off between Toffoli depth and clean ancilla, demonstrating that with additional ancilla, the exact Toffoli depth can be reduced to $2\log_{2}n$.
@@
 This step reduces the overall Toffoli count by $m_{1}$ compared to the circuit presented in khattar24 .
-The schematic diagram of the MCT decomposition using $m=m_{1}+m_{2}$ many ancilla is similar to the diagram we used for the exact Toffoli depth enumeration in Section IV.1.
+The schematic diagram of the MCT decomposition using $m=m_{1}+m_{2}$ many ancillae is similar to the diagram we used for the exact Toffoli depth enumeration in Section IV.1.
 $2m_{1}\rightarrow$ $m_{1}$
@@
 Thus, it can be observed that beyond a certain point, further increases in the number of ancilla qubits do not necessarily lead to a reduction in the Toffoli depth when using conditionally clean ancilla.
-The Toffoli depth of the MCT decomposition using $m=m_{1}+m_{2}$ many ancilla, can be estimated from the following lemma.
+The Toffoli depth of the MCT decomposition using $m=m_{1}+m_{2}$ many ancillae can be estimated from the following lemma.
 Lemma 4.
-The exact Toffoli depth of the Step II or Step IV of an $n$-MCT circuit decomposition, implemented above, using $m=m_{1}+m_{2}$ clean ancilla following the conditionally clean ancilla technique, varies with $n$ as follows.
+The exact Toffoli depth of Step II or Step IV of an $n$-MCT circuit decomposition, implemented above, using $m=m_{1}+m_{2}$ clean ancilla following the conditionally clean ancilla technique, varies with $n$ as follows.
 $\delta=\begin{cases}\lfloor\log_{2}m_{1}\rfloor+2k-3,\\ \indent\text{for }n\in\left[\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k-1}+k-1,\right.\\ \hskip 147.95424pt\left.m_{1}2^{k}+k-2\right]\\ \lfloor\log_{2}m_{1}\rfloor+2k-2,\\ \indent\text{for }n\in\left[m_{1}2^{k}+k-1,\right.\\ \hskip 73.97733pt\left.\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k}+k-1\right]\end{cases}$
@@
 Lemma 5.
-The number of control qubits, $\sigma$, in the Step III of an $n$-MCT circuit decomposition, implemented above using $m=m_{1}+m_{2}$ ancilla qubit, varies with $n$ as follows.
+The number of control qubits, $\sigma$, in Step III of an $n$-MCT circuit decomposition, implemented above using $m=m_{1}+m_{2}$ ancilla qubit, varies with $n$ as follows.
 $\sigma=\left\{\begin{array}[]{l}k+m_{1}-1,\\ \indent\text{ for }n\in\left[\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k-1}+k-1,\right.\\ \hskip 147.95424pt\left.m_{1}2^{k}+k-1\right]\\ k+m_{1},\\ \indent\text{ for }n\in\left(m_{1}2^{k}+k-1,\right.\\ \hskip 73.97733pt\left.\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k}+k-1\right]\end{array}\right.$
@@
 However, a significant part of the smaller MCT is applied simultaneously with Step II and Step IV, making the effective Toffoli depth of Step III, either $3$ or $5$.
-From above lemma, we can now estimate the exact Toffoli depth of the $n$-MCT circuit decomposition using $m=m_{1}+m_{2}$ ancilla qubits, as presented above.
+From the above lemma, we can now estimate the exact Toffoli depth of the $n$-MCT circuit decomposition using $m=m_{1}+m_{2}$ ancilla qubits, as presented above.
 Theorem 1.
@@
 It is now understood that the depth of $n$-MCT can be achieved in $\mathcal{O}(\log_{2}n)$, which can be explicitly written $c\log_{2}n$, where $c$ needs to be properly estimated, for exact trade-offs which is the main motivation in this paper.
-In Proposition 3, we have shown that in the MCT decomposition using two clean ancilla with the technique from khattar24 , $c$ is strictly greater than 3, in fact, it is around 4.
+In Proposition 3, we have shown that in the MCT decomposition using two clean ancillae with the technique from khattar24 , $c$ is strictly greater than 3; in fact, it is around 4.
 However, using the ancilla - Toffoli depth trade-off, $c$ can be further reduced to a certain extent.
@@
 From Proposition 1, it is evident that the Toffoli depth remains constant over a range of values for $n$.
-Since, here we are interested in showing the lower bound, we consider $n$ assumes the highest value in the range, and still show that $\log_{2}n$ is strictly less than the corresponding depth, as follows.
+Since here, we are interested in showing the lower bound, we consider $n$ assumes the highest value in the range and still show that $\log_{2}n$ is strictly less than the corresponding depth, as follows.
 Let us first proof the item 1. As the control states are not required to be returned to their original state, we consider the Toffoli depth due to Step I, Step II, and half of Step III only, which is
@@
 The above theorem shows that using the conditionally clean ancilla, the constant factor $c$ can never be reduced to $1$.
-In the next section, we demonstrate that the exact Toffoli depth in the Clifford + Toffoli decomposition of an $n$-MCT, is lower bounded by $\lceil\log_{2}n\rceil$, regardless of the technique or the number of ancilla qubits used.
+In the next section, we demonstrate that the exact Toffoli depth in the Clifford + Toffoli decomposition of an $n$-MCT is lower bounded by $\lceil\log_{2}n\rceil$, regardless of the technique or the number of ancilla qubits used.
 This bound is achievable through complete binary tree decomposition of $n$-controlled Toffoli gates.
@@
 To implement an $n$-MCT, we begin with $n$ leaf nodes, where each pair of control qubits is combined using a Toffoli gate, reducing the number of active qubits in each round.
-At each stage, either two parent nodes or one parent node along with a leftover node from the previous round are further combined into a new parent node.
+At each stage, either two parent nodes or one parent node, along with a leftover node from the previous round, are further combined into a new parent node.
 This process continues iteratively until the final Toffoli gate transfers the accumulated information to the root node (the target qubit).
@@
 ∎
-A complete binary tree structure, having $n$ leaf nodes, has a hight $\log_{2}n$, i.e., when $n$ is not a power of $2$.
+A complete binary tree structure, having $n$ leaf nodes, has a height $\log_{2}n$, i.e., when $n$ is not a power of $2$.
 When $n=2^{k}$, for some $k\in\mathbb{N}$, then it becomes a perfect binary tree, having the depth exactly $\log_{2}n=k$.
-In both the cases, the number of internal nodes, i.e., ancilla qubit is $n-2$.
+In both cases, the number of internal nodes, i.e., ancilla qubit, is $n-2$.
 Additionally, the Toffoli count for both these cases becomes $n-1$.
@@
 In the diagram, the first four Toffoli gates in the first two columns are implemented simultaneously, having depth 1.
-The next two Toffoli gates are applied sequentially, having Toffoli depth 1, each.
+The next two Toffoli gates are applied sequentially, having a Toffoli depth of 1 each.
 Figure 12: $7$-MCT circuit decomposition with a Toffoli depth 3.
@@
 Towards, the conclusion, let us outline a generalized approach that may provide a theoretical understanding.
-Given the Algebraic Normal Form (ANF) of a Boolean function, $f:\{0,1\}^{n}\rightarrow\{0,1\}^{m}$, naturally the maximum algebraid degree can be $n$.
-Thus, the reversible quantum circuit implementing $f$, can be designed with a Toffoli depth of $\lceil\log_{2}n\rceil$, utilizing an exponential number of ancilla qubits and $\mathcal{O}(n)$ Clifford gates.
+Given the Algebraic Normal Form (ANF) of a Boolean function, $f:\{0,1\}^{n}\rightarrow\{0,1\}^{m}$, naturally, the maximum algebraic degree can be $n$.
+Thus, the reversible quantum circuit implementing $f$ can be designed with a Toffoli depth of $\lceil\log_{2}n\rceil$, utilizing an exponential number of ancilla qubits and $\mathcal{O}(n)$ Clifford gates.
 This is because, in the ANF, the maximum degree is $n$, which means that $n$ different inputs are ANDed.
 Thus, an $n$-MCT is enough.
-Further, the ANF may contain at most $2^{n}$ terms, and that may require exponential number of ancilla qubits.
+Further, the ANF may contain at most $2^{n}$ terms, and that may require an exponential number of ancilla qubits.
 This is achieved by implementing all the required multi-controlled Toffoli (MCT) gates in parallel to compute the non-linear terms of the ANF.
@@
 In this direction, it is understood that the depth may really be quite less for each module to be implemented.
-Regarding the anicillae, for practical purposes, the ANF contains poly$(n)$ many terms in it, and in such cases the number of ancillae will also be poly$(n)$, instead of the generic exponential bound.
+Regarding the ancilla, for practical purposes, the ANF contains poly$(n)$ many terms in it, and in such cases, the number of ancilla will also be poly$(n)$ instead of the generic exponential bound.
 VI Conclusion
@@
 Finally, by incorporating Soeken’s measurement-based uncomputation technique (as referred in jaques19 ) for Toffoli decomposition, we extended this lower bound to T-depth as well.
+Acknowledgments: The authors like to thank Dr.
+Alessandro Luongo (CQT, NUS, and Inveriant Pte.
+Ltd., Singapore) for his valuable inputs while preparing this paper.
+Subhamoy Maitra acknowledges the funding support provided by the “Information Security Education and Awareness (ISEA) Project phase - III, Cluster - Cryptography, initiatives of MeitY, Grant No.
+F.No.
+L-14017/1/2022-HRD”.

### v3

@@
 (2024).
-In terms of a negative result, we first show that using such conditionally clean ancilla techniques, Toffoli depth can never achieve exactly $\lceil\log_{2}n\rceil$, though it remains of the same order.
-This highlights the limitation of the techniques exploiting conditionally clean ancilla [Nie et al., 2024, Khattar et al., 2024].
+In terms of a negative result, we first show that using such conditionally clean ancillae techniques, Toffoli depth can never achieve exactly $\lceil\log_{2}n\rceil$, though it remains of the same order.
+This highlights the limitation of the techniques exploiting conditionally clean ancillae [Nie et al., 2024, Khattar et al., 2024].
 Then we prove that, in a more general setup, the T-Depth in the Clifford + T decomposition, via Toffoli gates, is lower bounded by $\lceil\log_{2}n\rceil$, and this bound is achieved following the complete binary tree structure.
@@
 Unlike classical gates, quantum gates are inherently reversible and are mathematically represented by unitary matrices.
-The doubly-controlled X-gate, commonly known as the Toffoli gate, is among the most significant quantum gates, with critical applications in arithmetic operations, reversible computing, and oracle constructions wang2024boosting ; wang2024optimal ; wang2025comprehensive ; wang2024Minimum ; wang2024Reducing ; wang2023higher .
+The doubly-controlled X-gate, commonly known as the Toffoli gate, is among the most significant quantum gates, with critical applications in arithmetic operations wang2024boosting ; wang2024optimal ; wang2024Minimum ; wang2023higher , reversible computing lighter2019 ; dorcis2023 , and oracle constructions aes16 ; chacha21 ; fsr21 ; zuc24 .
 However, the Toffoli gate’s high resource demands, particularly in terms of T-Count, T-Depth, and ancilla qubits, can greatly influence the efficiency of fault-tolerant quantum circuits.
@@
 Over the past decades, numerous efforts have been made to reduce the resource requirements for the optimized implementation of multi-controlled Toffoli (MCT) gates miller11 ; saeedi13 ; maslov16 ; baker19 ; paler22 ; bala22 ; claudon24 .
-More recent works nie24 ; khattar24 have introduced the conditionally clean ancilla technique, which significantly reduces both the Toffoli depth and ancilla count in the MCT decomposition.
-Before proceeding further, let us explain the concept of conditionally clean ancilla that will be repeatedly referred to in this work.
+More recent works nie24 ; khattar24 have introduced the conditionally clean ancillae technique, which significantly reduces both the Toffoli depth and ancilla count in the MCT decomposition.
+Before proceeding further, let us explain the concept of conditionally clean ancillae that will be repeatedly referred to in this work.
 Generally, ancilla qubits are of two types, namely the clean ancilla, which is initialized to |$\ket{0}$ at the beginning of the circuit, and the dirty ancilla, which is initialized to some unknown quantum state at the beginning.
 The standard practice is to re-initialize the ancilla qubit to its initial state at the end of the computation.
-Conditionally clean ancilla nie24 ; khattar24 is the set of working (control) qubits, derived from the existing ones, that are treated as the clean ancilla based on certain assumptions, and re-initialized at the end of the computation.
-Although earlier works nie24 ; khattar24 did consider the dirty ancilla, in this paper, we are focusing on the conditionally clean ancilla derived from the clean ones only.
-Very recently, the conditionally clean ancilla technique has gained significant attention, including its applications in quantum adders remaud25 .
+Conditionally clean ancillae nie24 ; khattar24 are the set of working (control) qubits, derived from the existing ones, that are treated as the clean ancilla based on certain assumptions, and re-initialized at the end of the computation.
+Although earlier works nie24 ; khattar24 did consider the dirty ancilla, in this paper, we are focusing on the conditionally clean ancillae derived from the clean ones only.
+Very recently, the conditionally clean ancillae technique has gained significant attention, including its applications in quantum adders remaud25 .
 The resource optimization metrics under consideration are the Toffoli count, Toffoli depth, and Ancilla count.
 The Toffoli count and Toffoli depth are further refined with the T-Count and T-Depth as one may refer to Table 1.
-In this paper, we explore how the Toffoli depth (and consequently, T-Depth) can be reduced by increasing the number of clean ancilla qubits, utilizing the concept of conditionally clean ancilla.
+In this paper, we explore how the Toffoli depth (and consequently, T-Depth) can be reduced by increasing the number of clean ancilla qubits, utilizing the concept of conditionally clean ancillae.
 We also establish the limitation of this method in terms of the lower bound of the Toffoli depth.
@@
 In section IV, we explore the trade-off between the (clean) ancilla and the Toffoli depth using the existing techniques related to conditionally clean ones.
-In Section IV.1, we take a different look at viewing the MCT circuit decomposition using the conditionally clean ancilla technique of khattar24 and enumerate their exact Toffoli count.
-Following the trade-off, we show the reduction in Toffoli depth, and therefore in T-Depth (Construction 1, Example 1) compared to the recent work of Khattar and Gidney khattar24 , by introducing additional clean ancilla into the circuit, while keeping the Toffoli count constant.
+In Section IV.1, we take a different look at viewing the MCT circuit decomposition using the conditionally clean ancillae technique of khattar24 and enumerate their exact Toffoli count.
+Following the trade-off, we show the reduction in Toffoli depth, and therefore in T-Depth (Construction 1, Example 1) compared to the recent work of Khattar and Gidney khattar24 , by introducing additional clean ancillae into the circuit, while keeping the Toffoli count constant.
 These results are shown in Section IV.2.
@@
 In fact, this lower bound can be achieved for T-Depth as well, by the construction of jaques19 .
-That is, using the technique of jaques19 , we can obtain an $n$-MCT by further decomposing into the Clifford + T gate set with an exact T-Depth of $\lceil\log_{2}n\rceil$ too, using $22n-$ 2 ancilla, and a T-Count of $4(n-1)$.
+That is, using the technique of jaques19 , we can obtain an $n$-MCT by further decomposing into the Clifford + T gate set with an exact T-Depth of $\lceil\log_{2}n\rceil$ too, using $22n-$ 2 ancillae, and a T-Count of $4(n-1)$.
 Moreover, using the logical-AND circuit by Gidney gidney18 , the ancilla count can be reduced to $n-2$ keeping the T-Count constant.
@@
 In quantum computing, the Multi-Controlled Toffoli (MCT) gates play a crucial role in the design of complex quantum algorithms, including error correction codes and arithmetic operations.
-Despite its utility, the implementation of MCT gates exploiting the Clifford + T gate set requires careful decomposition to optimize resource usage, such as minimizing the T-Count, T-Depth, and the number of ancillary qubits.
+Despite its utility, the implementation of MCT gates exploiting the Clifford + T gate set requires careful decomposition to optimize resource usage, such as minimizing the T-Count, T-Depth, and the number of ancilla qubits.
 These optimizations are essential for practical quantum computation, where resource efficiency is a critical consideration.
 In this regard, we present a comprehensive overview of the existing benchmarks for Clifford + T decompositions of single and multi-controlled Toffoli gates in the following section.
-Let us now describe the idea of conditionally clean ancilla from khattar24 , as described in Figure 2.
+Let us now describe the idea of conditionally clean ancillae from khattar24 , as described in Figure 2.
 Given a clean ancilla, a Toffoli gate is implemented targeting the clean ancilla.
-If the clean ancilla is reversed, it means both the control qubits were |$\ket{1}$; thus, applying X gates on the control qubits will change them to |$\ket{0}$ and consequently can be used as conditionally clean ancilla in the next rounds.
-Similarly, in the following round, additional Toffoli gates are implemented targeting these conditionally clean ancilla, and applying the X gate on the control qubits will again make them conditionally clean for the next round, and the process continues until we exhaust all the control qubits.
+If the clean ancilla is reversed, it means both the control qubits were |$\ket{1}$; thus, applying X gates on the control qubits will change them to |$\ket{0}$ and consequently can be used as conditionally clean ancillae in the next rounds.
+Similarly, in the following round, additional Toffoli gates are implemented targeting these conditionally clean ancillae, and applying the X gate on the control qubits will again make them conditionally clean for the next round, and the process continues until we exhaust all the control qubits.
 Once the information from all the control qubits is accumulated in a few qubits, another (smaller) MCT gate is implemented with these qubits along with the ancilla as the control qubits and the original target as its target, so that if the ancilla was not modified in the first step, the target will also not be modified.
-Figure 2: Understanding the conditionally clean ancilla technique from khattar24 .
+Figure 2: Understanding the conditionally clean ancillae technique from khattar24 .
 Let us now analyze Figure 2 by distributing the possible scenarios in two mutually exclusive and exhaustive cases:
@@
 If one of $C_{0}$ or $C_{1}$ is |$\ket{0}$, then ancilla =|$\ket{0}$, and the target will not be flipped in Step III.
-If $C_{0}=C_{1}=\ket{1}$, but one of $C_{2},C_{3},C_{4}$ is |$\ket{0}$, then ancilla becomes |$\ket{1}$, and after applying X-gate on $C_{0},C_{1}$ will make, $C_{0}=C_{1}=\ket{0}$.
+If $C_{0}=C_{1}=\ket{1}$, but one of $C_{2},C_{3},C_{4}$ is |$\ket{0}$, then the ancilla becomes |$\ket{1}$, and after applying X-gate on $C_{0},C_{1}$ will make, $C_{0}=C_{1}=\ket{0}$.
 If one of $C_{2},C_{3}$ is |$\ket{0}$, then $C_{0}$ and consequently, $C_{1}$ also remains at |$\ket{0}$ state.
 Similarly, if $C_{4}=\ket{0}$, then also $C_{1}$ remains |$\ket{0}$, and in both scenario, the target will not be flipped in Step III.
-In Section IV.2, we modify the design by introducing additional ancilla qubits in Step I and subsequently creating more conditionally clean ancilla to begin with, which eventually reduces the overall Toffoli depth of the complete circuit.
+In Section IV.2, we modify the design by introducing additional ancilla qubits in Step I and subsequently creating more conditionally clean ancillae to begin with, which eventually reduces the overall Toffoli depth of the complete circuit.
 III Warm-up: A consolidated view on recent developments in Toffoli decomposition
@@
 In the same paper, the authors applied a meet-in-the-middle algorithm to present a Toffoli decomposition (Figure 3(c)) with a T-Depth of $3$ and an overall depth of $9$.
-For an exact decomposition of a single Toffoli gate (without measurement-based feedback), this is the lowest T-Depth that one can achieve without using any ancillary qubit.
+For an exact decomposition of a single Toffoli gate (without measurement-based feedback), this is the lowest T-Depth that one can achieve without using any ancilla qubit.
 Using one ancilla qubit, the authors also proposed a Toffoli decomposition circuit (Figure 4(a)) using $7$ T gates and $12$ Clifford gates, achieving a T-Depth $2$.
@@
 Since our primary focus here is to reduce Toffoli depth while increasing the clean ancilla count, we ignore the zero-ancilla implementation here.
-Figure 10: MCT decomposition using conditionally clean ancilla due to nie24 .
+Figure 10: MCT decomposition using conditionally clean ancillae due to nie24 .
 Proposition 2.
@@
 Since none of the Toffoli gates are applied simultaneously, the resulting Toffoli depth is $32n-$ 3, while the T-Depth is $32n-$ 3, following the T-Depth $1$ Toffoli implementation by jaques19 , requiring one more reusable ancilla.
-Furthermore, when the availability of clean ancilla increases to $2$ (Figure 11(b)), the Toffoli depth reduces to $\mathcal{O}(\log_{2}n)$ while maintaining the Toffoli count constant.
-Figure 11 illustrates the circuit diagrams of $10$-MCT using $1$ and $2$ clean ancilla, due to khattar24 .
-Additionally, if the clean ancilla is replaced with the dirty ancilla, the Toffoli count increases to $3216n-$ 32, and the Toffoli depth doubles under both scenarios.
-As we focus here on the conditionally clean ancilla derived solely from clean ancilla, we do not delve into an exact analysis of the Toffoli depth for the dirty ancilla circuit implementation.
+Furthermore, when the availability of clean ancilla qubit increases to $2$ (Figure 11(b)), the Toffoli depth reduces to $\mathcal{O}(\log_{2}n)$ while maintaining the Toffoli count constant.
+Figure 11 illustrates the circuit diagrams of $10$-MCT using $1$ and $2$ clean ancillae, due to khattar24 .
+Additionally, if the clean ancillae is replaced with the dirty ancillae, the Toffoli count increases to $3216n-$ 32, and the Toffoli depth doubles under both scenarios.
+As we focus here on the conditionally clean ancillae derived solely from clean ancilla qubits, we do not delve into an exact analysis of the Toffoli depth for the dirty ancilla circuit implementation.
 (a) Ancilla: 1, Toffoli depth: 17.
@@
 In this section, we provided the exact enumerations of Toffoli count and Toffoli depth of MCT circuit decomposition across various state-of-the-art results.
-For the exact enumeration of the Toffoli depth (and T-Depth), following khattar24 , we propose a novel approach of viewing the MCT decomposition with conditionally clean ancilla, which is a direct extension of khattar24 , thereby presented separately in Proposition 3 of Section IV.1.
+For the exact enumeration of the Toffoli depth (and T-Depth), following khattar24 , we propose a novel approach of viewing the MCT decomposition with conditionally clean ancillae, which is a direct extension of khattar24 , thereby presented separately in Proposition 3 of Section IV.1.
 Later, the approach has also been used to demonstrate the trade-off between Toffoli depth and the clean ancilla count.
@@
 IV Further Reduction of Toffoli Depth
-In this section, we adopt the MCT decomposition technique using the conditionally clean ancilla, as introduced in nie24 ; khattar24 , and propose a trade-off between the Toffoli depth with the clean ancilla, showing improvement over the recent work of Khattar and Gidney khattar24 .
+In this section, we adopt the MCT decomposition technique using the conditionally clean ancillae, as introduced in nie24 ; khattar24 , and propose a trade-off between the Toffoli depth with the clean ancilla, showing improvement over the recent work of Khattar and Gidney khattar24 .
 The authors proposed the MCT circuit decompositions techniques, achieving a Toffoli depth of $\mathcal{O}(n)$ with a single clean ancilla and $\mathcal{O}(\log_{2}n)$ with two clean ancillae, where in both cases, the Toffoli count becomes $32n-$ 3.
@@
 Step I and Step V combined have a Toffoli depth of 1 (using logical-AND).
-In Step II, information from the remaining $n-2$ control qubits is accumulated into $k(\ll n)$ qubits using the conditionally cleaned ancilla technique.
+In Step II, information from the remaining $n-2$ control qubits is accumulated into $k(\ll n)$ qubits using the conditionally cleaned ancillae technique.
 Next, in Step III, a $(k+1)$-MCT is implemented using the unused ancilla, following the single-ancilla technique, having a Toffoli depth of $12(k+1)-3=2k-$ 1.
@@
 Additionally, the number of layers ($k$) determines the size of the MCT gate required for Step III.
-In the above example, Step I and Step II of a $32$-MCT decomposition have been shown using the conditionally cleaned ancilla technique.
+In the above example, Step I and Step II of a $32$-MCT decomposition have been shown using the conditionally cleaned ancillae technique.
 The Step I has a Toffoli depth of $1$, and Step II has a Toffoli depth of $7$.
@@
 Clearly, the uncomputation (Step IV) requires an additional Toffoli depth of $8$, i.e., the Toffoli depth of the complete circuit is $19$.
-Using this new approach, the $5$-MCT circuit using a conditionally clean ancilla technique, presented in Figure 2, can be seen as follows.
+Using this new approach, the $5$-MCT circuit using a conditionally clean ancillae technique, presented in Figure 2, can be seen as follows.
 $2\rightarrow$ 11 11
@@
 Consequently, the overall Toffoli depth of the circuit remains unchanged.
-In the direction of enumerating the exact Toffoli depth of an $n$-MCT decomposition due to khattar24 , using 2 clean ancilla, we have the following lemma.
+In the direction of enumerating the exact Toffoli depth of an $n$-MCT decomposition due to khattar24 , using 2 clean ancillae, we have the following lemma.
 Lemma 1.
@@
 Proposition 3.
-Following khattar24 , the exact Toffoli depth of an $n$-MCT decomposition using $2$ clean ancilla is lower bounded by $3\log_{2}n$.
+Following khattar24 , the exact Toffoli depth of an $n$-MCT decomposition using $2$ clean ancillae is lower bounded by $3\log_{2}n$.
 Proof.
@@
 Although the above proposition establishes that the exact Toffoli depth using two clean ancillae, as per khattar24 , is lower bounded by $3\log_{2}n$, this bound is not tight and is closer to $4\log_{2}n$.
-In the following section, we analyze the trade-off between Toffoli depth and clean ancilla, demonstrating that with additional ancilla, the exact Toffoli depth can be reduced to $2\log_{2}n$.
+In the following section, we analyze the trade-off between Toffoli depth and clean ancilla, demonstrating that with additional ancilla qubit, the exact Toffoli depth can be reduced to $2\log_{2}n$.
 IV.2 Exact Trade-Off between Toffoli Depth and Clean Ancillae
@@
 For $m=3$, we set $m_{1}=2$ and $m_{2}=1$.
-Furthermore, for $m\geq 4$, we assume $m_{2}=2$ to implement the smaller MCT in Step III using the $2$-clean ancilla technique described in khattar24 .
+Furthermore, for $m\geq 4$, we assume $m_{2}=2$ to implement the smaller MCT in Step III using the $2$-clean ancillae technique described in khattar24 .
 The circuit design is as follows.
@@
 We implement them using the logical-AND circuit to avoid resource requirements for uncomputation in the final step.
-  • In Step II, information from the remaining control qubits is accumulated into $k(<n)$ qubits using the conditionally cleaned ancilla.
-  • In Step III, a $(k+m_{1})$-controlled Toffoli gate is implemented using the remaining $m_{2}$ many ancilla, following the $2$-ancilla technique, and the target qubit is modified.
+  • In Step II, information from the remaining control qubits is accumulated into $k(<n)$ qubits using the conditionally cleaned ancillae.
+  • In Step III, a $(k+m_{1})$-controlled Toffoli gate is implemented using the remaining $m_{2}$ many ancillae, following the $2$-ancilla technique, and the target qubit is modified.
   • In Step IV, uncomputation is performed to return the control qubits to their original state. As earlier, the Toffoli count and Toffoli depth of Step IV are identical to those in Step II.
-  • Finally, in Step V, the first $m_{1}$ many Toffoli gates are uncomputed to clean up the ancilla, using logical-AND based uncomputation.
+  • Finally, in Step V, the first $m_{1}$ many Toffoli gates are uncomputed to clean up the ancillae, using logical-AND based uncomputation.
 This step reduces the overall Toffoli count by $m_{1}$ compared to the circuit presented in khattar24 .
@@
 where the Toffoli depth for Step I is $1$, and the Toffoli depth for Step II (and Step IV) are $5$ each.
-Additionally, in Step III, we need to implement a $6$-controlled Toffoli using $m_{2}=2$ ancilla, which requires an additional Toffoli depth of $3$.
+Additionally, in Step III, we need to implement a $6$-controlled Toffoli using $m_{2}=2$ ancillae, which requires an additional Toffoli depth of $3$.
 Consequently, the Clifford + Toffoli decomposition of the $32$-controlled Toffoli gate, utilizing $(3+2)=5$ ancilla qubits, achieves a total Toffoli depth of $1+2(5)+3=14$.
@@
 where the Toffoli depth remains $14$.
-Thus, it can be observed that beyond a certain point, further increases in the number of ancilla qubits do not necessarily lead to a reduction in the Toffoli depth when using conditionally clean ancilla.
+Thus, it can be observed that beyond a certain point, further increases in the number of ancilla qubits do not necessarily lead to a reduction in the Toffoli depth when using conditionally clean ancillae.
 The Toffoli depth of the MCT decomposition using $m=m_{1}+m_{2}$ many ancillae can be estimated from the following lemma.
 Lemma 4.
-The exact Toffoli depth of Step II or Step IV of an $n$-MCT circuit decomposition, implemented above, using $m=m_{1}+m_{2}$ clean ancilla following the conditionally clean ancilla technique, varies with $n$ as follows.
+The exact Toffoli depth of Step II or Step IV of an $n$-MCT circuit decomposition, implemented above, using $m=m_{1}+m_{2}$ clean ancillae following the conditionally clean ancillae technique, varies with $n$ as follows.
 $\delta=\begin{cases}\lfloor\log_{2}m_{1}\rfloor+2k-3,\\ \indent\text{for }n\in\left[\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k-1}+k-1,\right.\\ \hskip 147.95424pt\left.m_{1}2^{k}+k-2\right]\\ \lfloor\log_{2}m_{1}\rfloor+2k-2,\\ \indent\text{for }n\in\left[m_{1}2^{k}+k-1,\right.\\ \hskip 73.97733pt\left.\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k}+k-1\right]\end{cases}$
@@
 Lemma 5.
-The number of control qubits, $\sigma$, in Step III of an $n$-MCT circuit decomposition, implemented above using $m=m_{1}+m_{2}$ ancilla qubit, varies with $n$ as follows.
+The number of control qubits, $\sigma$, in Step III of an $n$-MCT circuit decomposition, implemented above using $m=m_{1}+m_{2}$ ancilla qubits, varies with $n$ as follows.
 $\sigma=\left\{\begin{array}[]{l}k+m_{1}-1,\\ \indent\text{ for }n\in\left[\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k-1}+k-1,\right.\\ \hskip 147.95424pt\left.m_{1}2^{k}+k-1\right]\\ k+m_{1},\\ \indent\text{ for }n\in\left(m_{1}2^{k}+k-1,\right.\\ \hskip 73.97733pt\left.\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k}+k-1\right]\end{array}\right.$
@@
 Theorem 1.
-The exact Toffoli depth, $\delta$, of the complete $n$-MCT circuit decomposition using $m=m_{1}+m_{2}$ clean ancilla is given by
+The exact Toffoli depth, $\delta$, of the complete $n$-MCT circuit decomposition using $m=m_{1}+m_{2}$ clean ancillae is given by
 $\delta=\begin{cases}2\lfloor\log_{2}m_{1}\rfloor+4k,\\ \indent\text{for }n\in\left[\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k-1}+k-1,\right.\\ \hskip 147.95424pt\left.m_{1}2^{k}+k-2\right]\\ 2\lfloor\log_{2}m_{1}\rfloor+4k+2,\\ \indent\text{for }n\in\left[m_{1}2^{k}+k-1,\right.\\ \hskip 73.97733pt\left.\left(m_{1}+2^{\lfloor\log_{2}m_{1}\rfloor-1}\right)2^{k}+k-1\right]\end{cases}$
@@
 However, using the ancilla - Toffoli depth trade-off, $c$ can be further reduced to a certain extent.
-In the following subsection, we demonstrate that with the conditionally clean ancilla technique, $c$ always remains strictly greater than 1, regardless of the number of ancilla used.
-IV.3 Proving the Lower Bound on Toffoli Depth using Conditionally Clean Ancilla
-In this subsection, we show that the Toffoli depth of an $n$-MCT using the conditionally clean ancilla technique can never be reduced to $\lceil\log_{2}n\rceil$.
+In the following subsection, we demonstrate that with the conditionally clean ancillae technique, $c$ always remains strictly greater than 1, regardless of the number of ancilla qubits used.
+IV.3 Proving the Lower Bound on Toffoli Depth using Conditionally Clean Ancillae
+In this subsection, we show that the Toffoli depth of an $n$-MCT using the conditionally clean ancillae technique can never be reduced to $\lceil\log_{2}n\rceil$.
 Theorem 2.
@@
   1. 1.
-    Assuming that the control states do not need to be restored to their original values, using the conditionally clean ancilla technique, the exact Toffoli depth must be strictly greater than $\lceil\log_{2}n\rceil$, irrespective of the number of available ancilla.
+    Assuming that the control states do not need to be restored to their original values, using the conditionally clean ancillae technique, the exact Toffoli depth must be strictly greater than $\lceil\log_{2}n\rceil$, irrespective of the number of available ancilla.
   2. 2.
@@
 Therefore, Case II holds for all $k\geq 2$ for some $k\in\mathbb{N}$.
-In conclusion, using the conditionally clean ancilla technique, the exact Toffoli depth of an $n$-controlled Toffoli decomposition can not be reduced to $\lceil\log_{2}n\rceil$, even when the control qubits are not restored to their original states.
+In conclusion, using the conditionally clean ancillae technique, the exact Toffoli depth of an $n$-controlled Toffoli decomposition can not be reduced to $\lceil\log_{2}n\rceil$, even when the control qubits are not restored to their original states.
 Item 2, where the control qubits are restored to their original state, can be proved inductively in a similar manner.
@@
 ∎
-The above theorem shows that using the conditionally clean ancilla, the constant factor $c$ can never be reduced to $1$.
+The above theorem shows that using the conditionally clean ancillae, the constant factor $c$ can never be reduced to $1$.
 In the next section, we demonstrate that the exact Toffoli depth in the Clifford + Toffoli decomposition of an $n$-MCT is lower bounded by $\lceil\log_{2}n\rceil$, regardless of the technique or the number of ancilla qubits used.
@@
 By leveraging additional ancilla qubits, we achieved a lower Toffoli depth compared to their approach.
-Furthermore, we demonstrated that the conditionally clean ancilla technique cannot reduce the Toffoli depth strictly to $\lceil\log_{2}n\rceil$, irrespective of unlimited availability of clean ancilla.
+Furthermore, we demonstrated that the conditionally clean ancillae technique cannot reduce the Toffoli depth strictly to $\lceil\log_{2}n\rceil$, irrespective of unlimited availability of clean ancilla.
 Additionally, we established that, regardless of the decomposition technique or available ancillae, the Toffoli depth of an $n$-MCT circuit is fundamentally lower-bounded by $\lceil\log_{2}n\rceil$, with the optimal depth achieved through binary tree-based MCT decomposition.
@@
 Ltd., Singapore) for his valuable inputs while preparing this paper.
+Suman Dutta and Anupam Chattopadhyay acknowledge the support of ‘MoE AcRF Tier 1 award RT10/23’.
 Subhamoy Maitra acknowledges the funding support provided by the “Information Security Education and Awareness (ISEA) Project phase - III, Cluster - Cryptography, initiatives of MeitY, Grant No.